### v1

# PGD-based optimization of 3D bobsleigh track centerlines from 2D centerlines for simulation applications

Abstract
The centerline of a bobsleigh track defines its geometry and is essential for simulation modeling. To reduce bBobsleigh training costs, leveraging the centerline of the bobsleigh track to construct a virtual environment that closely replicates real competitive settings presents a promising solution. However, publicly available centerline data are typically limited and it is imprecise to construct a training system solely based on 2-dimensional (2D) centerline. To address this practical issue, this paper proposes a method for generating a 3-dimensional (3D) track centerline based on 2D centerline data. Incorporating international track design regulations, the method formulates an optimization problem that considers total track length, height difference, slope constraints, and geometric continuity. A Projected Gradient Descent (PGD) algorithm is used to solve the optimization problem. The generated 3D centerlines are compared with real track data, and the results show that the method can reproduce realistic centerline trends from original or scaled 2D data. For the selected track segment, the relative errors in total length, height difference, and average slope are within 1.7%, 3.2% and 4.1%, respectively, for real 2D data and within 1.1%, 3.5% and 4.3% respectively for scaled data. All slope values remain within the allowable limits. Moreover, by adjusting the segmentation or modifying the weight of height difference in the cost function, various centerline styles applicable to different competitions can be generated. Under different segmentation and weight factors, the maximum errors reach up to 4.4%, 4.8%, and 9.8%, and 4.4%, 4.8%, and 10.0%, respectively. The proposed method provides a flexible and efficient tool for supporting bobsleigh track centerline design.
keywords:
Bobsleigh, track centerline generation, international Olympic bobsleigh competition rules, projected gradient descent
1 Introduction
Since bobsleigh was formally recognized as a Winter Olympics event in 1924, it has attracted widespread attention from participating countries bi:bu4 . Bobsleighs are unpowered vehicles that rely entirely on the initial push from athletes and the conversion of gravitational potential energy into kinetic energy as they descend the track bi:1 ; 3 ; 5 . The competition is timed to an accuracy of milliseconds, and the final times often differ by only a few tenths of a second4 . To minimize energy loss, athletes typically avoid using the brake unless necessary before crossing the finish line. The dynamics of the bobsleigh and the driver’s control skill are key factors influencing performance bi:bu1 ; r2 .
However, due to restrictions from official governing bodies and the high cost of maintaining bobsleigh tracks7 , even countries with access to tracks often do not allow athletes to train on them regularly. Typically, limited test runs are permitted only shortly before official competitions bi:2 ; 1 ; bi:bu2 ; 6 . Consequently, simulation-based training and performance analysis have become vital tools for optimizing bobsleigh design and improving athlete technique. Researchers have developed virtual tracks, bobsleigh-blade/ice interaction models, multibody dynamics models of bobsleighs, and driver models. These simulation frameworks are increasingly being used to support bobsleigh structural design and help athletes devise optimal racing strategies. And there are limited studies on optimizing the centerline of bobsleigh tracks.
Early bobsleigh simulation research was significantly constrained by the limited computational capabilities of CPUs at the time, often relying on simplified point-mass models moving on smooth hyperbolic surfaces bi:3 ; bi:4 . The study bi:4 developed a partial surface model of a bobsleigh track, treating the bobsleigh as a point mass to analyze its dynamic response under external forces. However, the authors also acknowledged that due to the lack of available data, virtual training environments differ significantly from the real settings, limiting the variety of technical strategies athletes can efficiently practice. Based on this point-mass framework and real data from the Lillehammer track in Norway, The study bi:5 further designed an optimal controller to simulate bobsleigh driving behavior, aiding in athlete training. To better capture the complex nonlinear dynamics of bobsleigh motion on narrow, twisted, banked, and icy tracks, a geometric model is proposed bi:6 , which represents the tracks as a sequence of discrete points arranged along a 3D spatial curve. Moreover, a comprehensive simulation framework incorporating a blade–ice contact friction model is designed and validated on the Cesana and Whistler tracks, offering valuable guidance for track design and athlete training. The study bi:8 further validates its agreement between the simulated and experimental results on the Cesana track. However, the study did not involve real-time simulation, and the applicability of the proposed model to other official tracks worldwide was not examined. Subsequently, the geometric model is further extended into a multibody dynamics model in bi:2 , where more detailed dynamic characteristics are considered, resulting in better driving fidelity. Recent advances in bobsleigh simulation have focused on improving blade–ice friction modeling and virtual training environments bi:bu3 ; bi:7 . The study bi:10 developed longitudinal and lateral friction models through experiments, FEM and nonlinear regression, validating their effectiveness in simulators to evaluate driving styles. Some studies further integrate steering and visual feedback or design immersive systems for training and tourism bi:3 ; bi:2 ; bi:10 ; bi:11 . The study bi:11 introduced a VR hammock system to mimic sliding sensations, while the studies bi:12 ; bi:13 proposed automated methods for generating 3D track surfaces.
However, most of the aforementioned studies focus on the simulation of only one or two tracks, and the conclusions drawn lack generalizability. Many researchers have acknowledged the necessity of validating their proposed models against simulation results from other tracks in order to derive more universally applicable insights for athlete training. Currently, track data are primarily obtained through a combination of field measurements and industrial software modeling, which involves significant labor and material costs. Moreover, most countries do not publicly disclose the official data of their bobsleigh tracks. Therefore, acquiring a greater variety of available official track data worldwide has become an urgent need for enhancing both the diversity of research findings in simulation analysis and the general applicability of virtual training outcomes for bobsleigh athletes. The centerline of a bobsleigh track serves as a crucial dataset that defines the trajectory of the track. In contrast, the cross-sectional shapes are relatively standardized and can be generated using industrial software based on common templates bi:12 .
This study focuses on typical bobsleigh track centerlines that determine the trajectory characteristics of the track and proposes a method for generating 3D centerlines from 2D centerlines. In accordance with international track design regulations, an optimization problem is formulated that considers factors such as total track length, height difference, slope, and continuity. A PGD is used to solve the problem. The generated 3D centerline is quantitatively compared with actual track data, and the influence of different track characteristics and algorithmic parameters on the generation results is investigated. The proposed method provides data support for enhancing the diversity of theoretical research and improving the generalizability of training outcomes for bobsleigh athletes.
2 Methodology and Computational Framework
2.1 Formulation of the Optimization Problem
According to the International Olympic Bobsleigh Competition Rules bi:14 , a bobsleigh track consists of a sequence of straight and curved sections. These two types of track segments exhibit distinct geometric characteristics, which necessitate different approaches for their analysis and generation. Therefore, it is essential to segment the track into straight and curved sections. Since the curvature of a straight segment is zero, curvature serves as a key variable for distinguishing between straight and curved portions.
The horizontal plane is defined by the $x$ and $y$ axes, while the $z$-axis is oriented perpendicular to this plane. Taking the starting point of the track as the origin, a set of equally spaced points are sampled along the 2D centerline. The generation is shown in Figure 1. The 2D curvature at each point is then calculated by
$K_{pk}=\frac{|\ddot{x}\dot{y}-\dot{x}\ddot{y}|}{\left(\dot{x}^{2}+\dot{y}^{2}\right)^{3/2}}$ (1)
where $K_{pk}$ denotes the curvature at the current point, and $x$ and $y$ represent the coordinates of the point on the $x$-and $y$-axes, respectively. $\dot{x}$ and $\dot{y}$ are the first-order finite differences of the $x$ and $y$ coordinates at the current point, while $\ddot{x}$ and $\ddot{y}$ represent the corresponding second-order finite differences.
Figure 1: Transition from a 2D centerline to a 3D centerline.
According to the calculated curvature values, points with curvature below a predefined threshold are classified as points on straight segments. However, due to potential measurement errors in the raw data, individual points may be misclassified. To address this issue, the classification of each point is determined based on the classification of its neighboring region. Specifically, if the neighboring points are identified as part of a straight segment and the curvature at the current point is also below the threshold, the point is considered to lie on a straight segment. Otherwise, it is classified as part of a curved segment.
For bobsleigh tracks, slope and length are two critical characteristics. For a straight segment, the slope $g_{k}$ remains constant throughout. For a curved segment, when the distance $d$ between two points is extremely small, the segment between them can be approximated as a straight line. Therefore, if the coordinates of the previous point are $(x_{k-1},y_{k-1},z_{k-1})$, and the current point’s $x$ and $y$ coordinates are $(x_{k},y_{k})$, then the $z$-coordinate of the current point can be approximated as $z_{k}=z_{k}-d\cdot g_{k})$. The distance $d$ is calculated as follows:
$d=\sqrt{(x_{k}-x_{k-1})^{2}+(y_{k}-y_{k-1})^{2}}$ (2)
The starting point of the track is defined as $(0,0,H)$, where $H$ represents the height difference along the $z$-axis between the highest point at the start and the lowest point of the entire track. Therefore, it is necessary to determine the height and slope of each track segment, which allows the 3D coordinates of each point to be recursively computed.
The generated 3D coordinates should satisfy several constraints:
  1. 1.
    The slope of each segment must remain within a reasonable range;
  2. 2.
    The total length of the generated 3D track should be equal to the known total track length $L$;
  3. 3.
    The total height difference of the generated track must match the known value $H$;
  4. 4.
    To ensure the continuity of the bobsleigh track, the curvature in the $z$-axis direction should be significantly smaller than that in the $x$ and $y$ axes, meaning that the 2D curvature at each point should approximately equal the 3D curvature;
  5. 5.
    The average slope of the track should be equal to the known average slope $\overline{g_{k}}$.
In summary, suppose the 2D coordinates of the track consist of $n$ points $(x_{k},y_{k})$, where $k=1,2,\ldots,n$, and the 2D curvature at each point is denoted by $K_{pk}$. Let the track be divided into $j$ segments, with the slope of each segment denoted by $g_{m}$, where $m=1,2,\ldots,j$. The generated 3D coordinates for each point are $(x_{k},y_{k},z_{k})$, and the 3D curvature at each point is $K_{sk}$. Let $L$ be the actual total track length, $H$ the actual total height difference, and $\bar{g}$ the actual average slope.
The generation of the 3D track from the 2D centerline can then be formulated as an optimization problem, as shown in Equation (3), in which the slope of each segment is optimized such that the resulting 3D coordinates satisfy the five aforementioned constraints.
$J=\min_{g_{k}}\Bigg{[}$ $a\left(L-\sum_{k=2}^{n}d_{k}\right)+b\left(H-\sum_{k=2}^{n}\sqrt{(z_{k}-z_{k-1})^{2}}\right)$ (3)
$+\frac{c}{n}\sum_{k=1}^{n}(K_{pk}-K_{sk})+d\left(\frac{1}{j}\sum_{m=1}^{j}g_{m}-\bar{g}\right)\Bigg{]}$
[…]
In this equation, $J$ represents the cost function to be minimized, and $j$ is the number of track segments. The parameters $a$, $b$, $c$, and $d$ are the weighting factors corresponding to the total length difference, height difference, curvature difference between the 2D and 3D centerlines, and the average slope difference, respectively. $g_{k_{\text{max}}}$ and $g_{k_{\text{min}}}$ denote the maximum and minimum allowable slopes. $d_{k}$ is the 3D distance between point $k$ and point $k-1$, calculated as follows:
$d_{k}=\sqrt{(x_{k}-x_{k-1})^{2}+(y_{k}-y_{k-1})^{2}+(z_{k}-z_{k-1})^{2}}$ (4)
The calculation formula for the 3D curvature $K_{sk}$ at each point is as follows:
$K_{sk}=\frac{\sqrt{\lvert u\rvert^{2}\lvert v\rvert^{2}-(vu)^{2}}}{\lvert v\rvert^{3}}$ (5)
where $u=[\dot{x},\dot{y},\dot{z}]$ is the first derivative vector of the coordinates, and $v=[\ddot{x},\ddot{y},\ddot{z}]$ is the second derivative vector of the coordinates.
To solve the bobsleigh track optimization problem established, a PGD bi:15 is used for parameter optimization. Compared to more complex methods such as genetic algorithms or simulated annealing, gradient descent offers faster convergence and requires fewer hyperparameters, making it more practical for repeated evaluations during track generation.
In this study, the height difference $h_{m}$ of each track segment is a high-dimensional parameter (with dimensionality corresponding to the number of track segments), which introduces significant nonlinearity to the optimization problem, making the exact computation. Therefore, the variation of the cost function is employed as a substitute for gradient computation. This approach does not rely on the explicit calculation of partial derivatives but instead approximates the gradient through a numerical difference method.
More specifically, the direction of the gradient is approximated by evaluating the change in the cost function value between two successive iterations, rather than explicitly computing all partial derivatives. The underlying concept is similar to the finite difference method; however, it places greater emphasis on variations in the cost function itself rather than on exact gradients. The corresponding computation is given as follows:
$\frac{\partial J}{\partial h_{i}}\approx\frac{J(h_{i})-J(h_{i-1})}{h_{i}-h_{i-1}}$ (6)
where $i$ denotes the iteration index.
If the optimization is based on 2D coordinate data with uniform scaling, an additional optimization variable, the scaling factor $f$, must be introduced. In this case, the corresponding computation is modified as follows:
$\frac{\partial J}{\partial f_{i}}\approx\frac{J(f_{i})-J(f_{i-1})}{f_{i}-f_{i-1}}$ (7)
The gradient descent update rule is given as follows:
$h_{i+1}=h_{i}-\eta\frac{\partial J}{\partial h_{i}}$ (8)
where $\eta$ is the learning rate, which is set to $10^{-5}$.
If the scaling factor is considered, it should also be updated using gradient descent as follows:
$f_{i+1}=f_{i}-\eta\frac{\partial J}{\partial f_{i}}$ (9)
Compared to SGD(Stochastic Gradient Descent) bi:sgd1 ; bi:sgd2 ; bi:sgd3 ; bi:sgd4 , PGD can be regarded as projecting the height difference of each segment during optimization to ensure that the corresponding slope of each segment falls within the prescribed limits, which in turn satisfies the overall constraint conditions. Although the projection is not explicitly performed on each variable (e.g., the height difference of each segment) at every gradient descent step, it is effectively realized by adjusting the slope of each segment to meet the constraints.
During the gradient descent process, the height difference of the current segment is denoted as $h_{m}$, and the corresponding slope is calculated as:
$g_{m}=\frac{h_{m}}{d_{m}}$ (10)
where $d_{m}$ represents the total length of a specific segment of the track, and it is calculated as follows:
$d_{m}$ $=\sum_{k=p+p_{2}+\cdots+p_{m-1}}^{p+p_{2}+\cdots+p_{m}}d_{k}$
where $p_{m}$ denotes the number of points within segment $m$ of the track. The slope $g_{m}$ is constrained to remain within a predefined range. If $g_{m}$ exceeds this range (e.g., smaller than min slope or larger than max slope), the corresponding height difference $h_{m}$ is adjusted accordingly:
{$h_{m}&=g_{k_{\text{min}}}\times d_{m},\quad&&g_{m}\leq g_{k_{\text{min}}}\\ h_{m}&=g_{k_{\text{max}}}\times d_{m},\quad&&g_{m}\geq g_{k_{\text{max}}}\end{aligned}\right.$ (11)
$\Delta J=|J_{i+1}-J_{i}|<\delta$ (12)
where $\delta$ represents the threshold of variation, which is set to $10^{-3}$.
2.2 Numerical Procedure
Based on the optimization equation formulated in accordance with the track design rules and the PGD, the computational procedure is illustrated in the following figures. First, the initial segmental height differences and the corresponding cost function value are calculated. Then, using a randomly assigned initial gradient, the segmental height differences are updated, and the $z$-coordinates of all points are recalculated based on these updated values.
Subsequently, the PGD algorithm enters an iterative loop. At each iteration, it first checks whether the slope of each track segment falls within the specified bounds. If any segment’s slope exceeds the upper or lower limits, the projection method is applied to map it back into the feasible range. Using the projected slope values, the segmental height differences are updated, followed by recalculating the $z$-coordinates of all points.
The cost function is then evaluated again, and the height differences are further updated. This process repeats until the change in the cost function value converges to a predefined threshold. At that point, the final $z$ for all points on the track are determined.
If a scaling factor is included in the optimization, the $x$ and $y$ are scaled at the beginning of each iteration, and the scaling factor is updated at the end of each iteration. The optimization process is illustrated in Figure 2.
Figure 2: Calculation flow diagram.
3 Results Analysis: Model Verification
3.1 Track with Real 2D Plane Data
Since some studies include real 2D plane data for a limited number of tracks bi:6 , this data can be used for solving the optimization problem. The generated track data are then compared with the actual track data from the studies. In this section, the impact of the track’s ascent sections on the results is not considered. Therefore, only the data from the track’s descent sections are used for track generation and comparison analysis. In addition, the generation of the track centerline is a non-convex optimization problem, which may have multiple local optima. As a result, the algorithm may converge to different solutions in different runs. While the overall trend of the generated centerlines remains consistent, local variations may occur between runs. This characteristic enables the generation of diverse track designs, providing multiple alternative options for practical use and enhancing the flexibility and applicability of track design. Therefore, three random results are generated each time for comparing with the actual track centerline.
3.1.1 Whistler Track
The actual design geometric parameters of the Whistler track are shown in Table 1.
Table 1: Actual geometric parameters of the Whistler track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1278 149.3 20.4 11.6
For the Whistler track, Figures 3 and 4 show the comparison of the actual track centerline’s 2D curvature and the actual and generated track centerlines’ 3D curvatures as a function of track distance. From Figures 3 and 4, it can be observed that there is no significant difference between the 2D curvature of the actual track centerline and the 3D curvature of both the actual and generated tracks across the track distance. Only small differences are found near the extreme values. Therefore, the impact of this factor is not further analyzed in the results.
Figure 3: Comparison of the planar curvature and spatial curvature of the actual Whistler track centerline with track distance.
Figure 4: Comparison of the 2D curvature and 3D curvature of the actual bobsleigh track centerline with track distance.
Table 2 shows the comparison of the generated and actual Whistler bobsleigh track centerline geometric parameters. From Table 2, it can be observed that the generated track’s total length, height difference, and average slope are all close to those of the actual track, with maximum errors of 1.7%, 3.2%, and 4.1%, respectively. Additionally, the slope values do not exceed the defined limits.
Table 2: Comparison of geometric parameters of the generated and actual Whistler track centerlines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1291.1 150.9 No 0.1208
Generated track 2 1290.2 144.5 No 0.1148
Generated track 3 1290.7 148.4 No 0.1177
Actual track 1278 149.3 No 0.1160
Figure 5 shows the comparison of the track height variation along the distance for the generated and actual Whistler track centerlines. From Figure 5, it can be seen that the three generated track centerlines closely follow the trend of the actual track centerline, with the track height continuously decreasing as the track distance increases. Specifically, for Generated Track 1, the height is lower than the actual track up to approximately 400 m, but higher beyond this point. This is because the average slope of each segment beyond 400 m is higher than that of the actual track, resulting in a larger height difference. For Generated Track 3, the height closely matches the actual track up to approximately 100 m, then falls below the actual track between 100 m and 300 m, and exceeds the actual track between 300 m and 900 m, before aligning with the actual track again beyond 900 m. This is because the randomly assigned slopes and height differences are close to those of the actual track. Generated Track 2 is slightly lower than the actual track between 0–200 m and beyond 1025 m, but closely matches the actual track between 200 m and 1025 m. This is due to the randomly assigned slopes and height differences being lower than those of the actual track.
Figure 5: Comparison of the track height variation along the distance for the generated and actual Whistler track centerlines.
Figure 6 shows the comparison of the track slope variation along the distance for the generated and actual Whistler track centerlines. From Figure 6, it can be seen that according to the study bi:6 , at track distances of approximately 100m, the model calculated in this paper results in a centrifugal force that exceeds the maximum allowable centrifugal force. Although the track slope does not exceed the design slope value of 0.204, the higher average slope, shorter total length, and larger height difference contribute to the centrifugal force exceeding the maximum allowable value. The different generated track slopes in this algorithm exhibit consistent trends along the track distance, with only minor variations in amplitude. All generated track slopes are within the design value range, demonstrating that the algorithm can provide a rich variety of track generation results, and that relevant parameters can be constrained to meet the track design requirements.
Figure 6: Comparison of the track slope variation along the distance for the generated and actual Whistler track centerlines.
3.1.2 Censana Track
Similar to the results of Whistler track, the actual geometric parameters of the Cesana track are shown in Table 3.
Table 3: Actual geometric parameters of the Cesana track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1359.2 124.5 18.3 9.1
Table 4 shows the comparison of geometric parameters of the generated and actual Cesana track centerlines. From Table 4, the simulated values of the generated track’s total length, height difference, and average slope are all close to those of the actual track, with maximum errors of 0.81%, 1.2%, and 2.6%, respectively. Moreover, all slope values remain within the prescribed limits.
Table 4: The comparison of geometric parameters of the generated and actual Cesana track centerlines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1360.3 124.0 No 0.0934
Generated track 2 1359.0 123.1 No 0.0926
Generated track 3 1359.0 123.0 No 0.0930
Actual track 1359.2 124.5 Exceeds by approximately 10m 0.0910
Figure 7 shows the comparison between the track height along the centerline of the generated and actual Cesana tracks as a function of distance. From Figure 7, the three centerlines generated by the proposed algorithm all follow a similar trend to that of the actual track, with the track height decreasing as the distance along the track increases.
For the generated Track 1, the height is close to that of the actual track before approximately 600 m, higher than the actual track between approximately 600 m and 1000 m, and lower than the actual track beyond approximately 1000 m. This is because the randomly assigned slope values in the segment between approximately 600 m and 1000 m are on average greater than those of the actual track. The result for generated Track 2 is similar to that of Track 1, with an even better match to the actual track before approximately 600 m. For generated Track 3, the result is similar to that of Track 1 before approximately 600 m, closely aligned with the actual track between approximately 600 m and 1000 m, but shows noticeable deviation after approximately 1000 m. This is due to the relatively smaller slope values assigned between approximately 600 m and 1000 m and the slightly lower overall height compared to the actual track.
Figure 7: Comparison of track height variation along the distance for the generated and actual Cesana track centerlines.
Figure 8 shows the comparison of track gradient along the centerlines between the generated and actual Cesana track. As shown in Figure 8, although the maximum designed gradient value is 0.183 bi:6 , the actual value exceeds 0.2, which may lead to a centrifugal force exceeding the allowable limit of 5g during cornering 8 . According to the calculation model presented in the study bi:6 , the maximum centrifugal force is approximately 4.5g, which can be attributed to the moderate height difference and sufficiently long total track length. While a lower gradient is critical for safety, an overly small gradient may reduce the differentiation in athlete performance, resulting in smaller performance gaps. The gradients of the tracks generated by the proposed algorithm follow a similar trend along the track distance, with only local variations in magnitude. All gradient values remain within the design limits, demonstrating that the gradient profile of the generated centerlines meets the required specifications.
Figure 8: Comparison of the track slope variation along the distance for the generated and actual Cesana track centerlines.
In summary, the proposed optimization equation can generate track centerlines with diverse characteristics based on real 2D data, and the trend of the generated centerline’s characteristic parameters is consistent with that of the actual track. The generated centerlines effectively cover the spatial domain of the original track centerlines. Simulation results on two selected tracks show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks, with maximum deviations of 1.7%, 3.2%, and 4.1%, respectively. Moreover, the gradients remain within the specified design limits, thereby validating the effectiveness of the proposed track centerline generation algorithm. The generation results are primarily influenced by parameters such as the randomly assigned segment gradients, height difference, distance, average gradient, and maximum gradient. In addition, the two real tracks analyzed herein highlight potential directions for further optimization.
3.2 Tracks Proportionally Scaled from Real 2D Planar Data
As real bobsleigh tracks with complete 2D planar data are scarce in the studies, while proportionally scaled versions of such data are relatively more available, proportionally scaled tracks are used here to validate the proposed track centerline generation model. By manually scaling real track data, this study simulates the case where original 2D data are unavailable, but proportionally scaled data are accessible. With the introduction of a scaling factor, the optimization equation is solved accordingly, and the generated track centerlines are compared with those of the actual tracks in terms of characteristic parameters.
Taking the Whistler track as an example in this section, the optimization problem considers a scaling factor $f$, where the horizontal $x$ and $y$ coordinates of the Whistler track are uniformly scaled to 90% of their original values (isotropic scaling is applied, with no inconsistency between the scaling ratios of $x$ and $y$ coordinates).
Figure 9 presents a comparison between the generated and actual $x$ and yy coordinates of the track centerline. As shown in Figure 9, the $x$ and $y$ coordinates calculated by the proposed algorithm follow the same trend as the actual values. The computed scaling factors $f$ are 1.1102, 1.1196, and 1.0857, corresponding to relative errors of 0.1%, 7.6%, and 2.3%, respectively (with respect to $f$ = 1.1). These results demonstrate that the proposed algorithm can effectively estimate the original $x$ and $y$ coordinates of the track centerline prior to scaling.
Figure 9: Comparison of the track slope variation along the distance for the generated and actual Cesana track centerlines.
To further analyze the computational results, Table 5 presents a comparison between the geometric parameters of the generated and actual track centerlines. As shown in Table 5, the simulated values of track length, height difference, and average slope of the generated track are all close to those of the actual track, with maximum errors of 1.1%, 3.5%, and 4.3%, respectively. Additionally, all slope values remain within the prescribed limits.
Table 5: The comparison of geometric parameters of the generated and actual track centerlines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1291.2 146.9 No 0.1150
Generated track 2 1290.8 143.9 No 0.1126
Generated track 3 1292.2 154.6 No 0.1210
Actual track 1278 149.3 No 0.1160
Figure 10: Comparison of track height variation along the distance for the generated and actual track centerlines.
Figure 11: Comparison of the track slope variation along the distance for the generated and actual track centerlines.
Figures 10 and 11 respectively compare the variation of track height and slope along the distance between the generated and actual track centerlines. Similar to the results in model verification section, the height variation trends of the centerlines generated by the proposed algorithm are consistent with those of the actual track, and all slope values of the generated centerlines remain within the design limits.
In summary, the proposed algorithm can compute the original (pre-scaled) $x$ and $y$ coordinates from the scaled $x–y$ plane coordinates. Similar to the algorithm based on the original unscaled $x–y$ plane coordinates, the one that incorporates the scaling factor is also capable of generating track centerlines with diverse characteristics while maintaining consistency with the overall trends of the actual track centerline. The generated centerlines sufficiently cover the range of the original centerline. Validation using the selected track demonstrates that the generated track’s length, height difference, and average slope are all close to the actual values, with maximum errors of 1.1%, 3.5%, and 4.3%, respectively. Moreover, the slope values do not exceed the specified limits, thereby verifying the effectiveness of the proposed track generation algorithm.
The track generation results are mainly influenced by parameters such as the randomly assigned slope of each segment, the track’s height difference, length, average slope, and maximum slope.
4 Results Analysis: Influence of Different Parameters on Track Generation with Actual 2D Plane Data
In this section, the Whistler and Cesana tracks are selected as case studies to analyze the effects of segment count and weighting factors on track generation under the condition of actual 2D plane data.
4.1 Whistler Track
4.1.1 Influence of Track Segment Count
For the Whistler track, the effect of dividing the track into 15, 26, and 33 segments on the generation results was analyzed. Table 6 presents a comparison of the geometric parameters between the generated and actual Whistler track centerlines. As shown in Table 6, under different segmentation settings, the simulated values of track length, height difference, and average slope are all close to those of the actual track, with maximum errors of 1.1%, 4.8%, and 7.7%, respectively. Additionally, all slope values remain within the specified limits.
Figure 12 shows the comparison of height variation along the distance between the generated and actual track centerlines. It can be observed that the height profiles of the generated centerlines follow the same trend as that of the actual track. For the Whistler track, increasing the number of track segments results in greater local slope variation, but the overall variation remains relatively minor. This is due to the increased number of local slope changes caused by a higher segmentation count. Conversely, when the number of segments is lower, local variations are reduced, while the overall slope variation becomes more pronounced, leading to better agreement with the actual track.
In this case, as the track distance increases, the height also increases. In the 0–600 m segment of the track, the generated centerlines based on 15 and 33 segments show higher consistency with the actual track. In the 600–1200 m segment, the track centerline generated with 15 segments aligns more closely with the actual track. This is because in the 0–600 m segment, the actual centerline exhibits a relatively steep slope, so a larger slope value in the generated centerline leads to a closer match with the actual track. In contrast, in the 600–1200 m segment, the actual centerline has a gentler slope, so a smaller slope value in the generated centerline results in better agreement with the actual track.
Table 6: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=15$ 1291.5 151.4 No 0.1198
Track segment count $s=26$ 1291.7 156.5 No 0.1249
Track segment count $s=33$ 1291.7 156.2 No 0.1222
Actual track 1278 149.3 No 0.1160
Figure 12: Comparison of track height variation along the distance for the generated and actual track centerlines.
4.1.2 Influence of The Weighting Factor
The height difference largely determines the elevation change of each segment along the track. Therefore, this section analyzes the influence of the weighting factor for height difference on the centerline generation results. The number of track segments is set to 26. Table 7 presents the comparison of geometric parameters between the generated and actual track centerlines. As shown in Table, the simulated values of track length, height difference, and average gradient under different height difference weighting factors are all close to the actual values. The maximum errors are 0.7%, 1.7%, and 3.8%, respectively. Furthermore, all slope values remain within the specified limits. As the height difference weighting increases, the height difference of the generated centerline becomes more consistent with the actual value.
Figure 13 shows a comparison between the generated and actual centerline track height variation with distance. As illustrated in Figure 13, the variation trend of the generated centerline height is consistent with that of the actual track. For the segment between 200 m and 600 m, the generated centerlines with weighting factors of 0.7 and 0.9 demonstrate a higher degree of agreement with the actual data. For the segment between 600 m and 1000 m, those with weighting factors of 0.5 and 0.7 are more consistent with the actual centerline.
This is because for similar track lengths, smaller height differences generally correspond to gentler gradients. In the segment between 200 m and 600 m, the generated centerline with a larger gradient shows better agreement with the actual track, whereas in the segment from 600 m to 1000 m, the centerline with a smaller gradient corresponds more closely to the actual values.
Table 7: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.5$ 1287.6 151.9 No 0.1183
Weight factor $b=0.7$ 1287.6 151.4 No 0.1197
Weight factor $b=0.9$ 1287.5 151.1 No 0.1204
Actual track 1278 149.3 No 0.1160
Figure 13: Comparison of track height variation along the distance for the generated and actual track centerlines.
4.2 Cesana Track
4.2.1 Influence of Track Segment Count
For the Cesana track, the influence of dividing the track into 12, 21, and 29 segments on the centerline generation results is analyzed. Table 8 presents the comparison of geometric parameters between the generated and actual track centerlines. As shown in Table 8, under different numbers of segments, the simulated values of track length, height difference, and average gradient are all close to the actual values, with maximum errors of 0.2%, 0.9%, and 2.4%, respectively. In all cases, the gradient remains within the specified limits.
Figure 14 shows the variation of track height along the distance for both the generated and actual centerlines. As can be seen, the generated centerline height follows a trend consistent with the actual track. For the segments from 0 to 600 m and after 1000 m, the centerline generated with 29 segments exhibits the highest degree of agreement with the actual centerline. In contrast, for the segment from 600 m to 1000 m, the centerline generated with 12 segments shows the highest level of agreement.
This is because increasing the number of segments leads to more local variations in the track gradient, while the overall variation remains limited. Conversely, a smaller number of segments results in fewer local variations but more pronounced overall gradient changes. For this particular track, in the segments from 0 to 600 m and after 1000 m, where the actual centerline has relatively steep gradients, the generated centerline better matches the actual one when the gradient values are higher. In the segment from 600 m to 1000 m, where the actual track gradient is gentler, the generated centerline more closely matches the actual track when the gradient values are smaller.
Table 8: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=12$ 1356.6 123.4 No 0.0932
Track segment count $s=21$ 1356.4 124.0 No 0.0932
Track segment count $s=29$ 1356.6 123.9 No 0.0931
Actual track 1359.2 124.5 No 0.0910
Figure 14: Comparison of track height variation along the distance for the generated and actual track centerlines.
4.2.2 Influence of Weighting Factor
This section also analyzes the influence of the height difference weight factor on the results of centerline generation, with the number of track segments set to 12. Table 9 presents the comparison of geometric parameters between the generated and actual track centerlines. As shown in Table 9, under different weight factors for height difference, the generated track length, height difference, and average slope are all close to those of the actual track, with maximum errors of 8%, 9%, and 5%, respectively, and all slope values remain within the allowable range.
Figure 15 shows the comparison of track height versus distance between the generated and actual track centerlines. As seen in Figure 15, in the section of 0–600 m, the generated track centerline with a weight factor of 0.1 exhibits better agreement with the actual centerline. In the section of 600–1300 m, the generated centerlines with weight factors of 0.7 and 0.9 show better agreement with the actual track.
This is because for similar track distances, a smaller height difference generally results in a smoother profile as indicated by the average slope. In the 0–600 m section, a larger slope in the generated track centerline better matches the actual track, while in the 600–1300 m section, a smaller slope in the generated centerline aligns more closely with the actual track centerline.
Table 9: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.1$ 1357.1 128.9 No 0.0974
Weight factor $b=0.7$ 1356.8 125.0 No 0.0940
Weight factor $b=1.3$ 1356.7 124.7 No 0.0941
Actual track 1359.2 124.5 No 0.0910
Figure 15: Comparison of track height variation along the distance for the generated and actual track centerlines.
To summarize, by dividing the track into different numbers of segments or adjusting the weight of height difference in the optimization equation, multiple centerlines with distinct characteristics can be generated. Within the parameter ranges used in this study, the generated track length, height difference, and average slope under different segmentations are all close to those of the actual tracks, with maximum errors of 1.1%, 4.8%, and 7.7%, respectively. Under different height difference weights, the maximum errors in track length, height difference, and average slope are 0.7%, 3.4%, and 7.0%, respectively. Moreover, all generated slopes remain within the prescribed limits. The agreement between the generated and actual centerlines varies depending on the track and the segmentation or weight factor used. The variety of centerlines generated by the proposed algorithm can be flexibly applied to meet specific track design requirements, thereby serving as a valuable aid in centerline design.
5 Results Analysis: Influence of Different Parameters on Track Generation under Scaled 2D Planar Data
In this section, the Nagano track in Japan and the Pyeongchang track in South Korea are selected as case studies to analyze the influence of segmentation and weight factors on track generation based on scaled 2D planar data.
5.1 Nagano Track
This section selects the Nagano track and the PyeongChang track as case studies to analyze the influence of the number of segments and the weight coefficient on the generated track under conditions involving scaled 2D planar data. Since the starting point of the final ascending segment of the track cannot be determined and the slope design rules for the ascending segment are similar to those for the descending segment bi:14 , the final ascending segment is treated as part of the descending segment. Its height change is included in the total height difference for analysis.
Under the algorithm proposed in this study, an analysis was conducted on the proportionally scaled 2D data of the Nagano track. The actual geometric parameters of the track are listed in Tables 10 bi:14 .
Table 10: Actual geometric parameters of the Nagano track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1700 118 15 7
5.1.1 Track Generation
Under the proposed algorithm, a track generation analysis is conducted based on the uniformly scaled 2D data of the Nagano track in Japan. The calculation results are shown in Table 11. Similar to the results in model verification section, the generated track length, height difference, and average slope are all close to the actual values, with maximum errors of 4.3%, 4.8%, and 8.6%, respectively. Moreover, all slope values remain within the prescribed limits.
Table 11: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1625.8 123.7 No 0.0760
Generated track 2 1652.4 115.6 No 0.0699
Generated track 3 1658 113.9 No 0.0686
Actual track 1700 118 No 0.0700
Figure 16: Comparison of track height variation along the distance for the generated and actual track centerlines.
5.1.2 Influence of Track Segment Count
For the Nagano track, the influence of dividing the track into 10, 15, and 21 segments on the centerline generation results is analyzed. Table 12 presents a comparison of the geometric parameters between the generated and actual track centerlines. As shown Table 12, the track length, height difference, and average slope of the generated track under different numbers of track segments are all close to those of the actual track, with maximum errors of 4.4%, 3.6%, and 8.6%, respectively. Furthermore, all slope values remain within the specified limits.
Figure 18 shows the track height variation along the distance for the generated centerlines. Similar to the results of track generation with actual 2D plane data section, Figure 18 demonstrates that the height of the generated centerline varies with distance in accordance with the actual track. A greater number of track segments results in more variations in local slopes while the overall change remains relatively small. This is because increasing the number of segments leads to more pronounced local slope fluctuations. Conversely, with fewer segments, the local variation is less significant, but the overall slope changes more substantially.
Table 12: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=10$ 1653.3 113.8 No 0.0688
Track segment count $s=15$ 1625.8 123.7 No 0.076
Track segment count $s=21$ 1656.4 117.3 No 0.0708
Actual track 1700 118 No 0.07
Figure 17: Track height variation along the distance for the generated centerlines.
This section also analyzes the effect of the weight factor for height difference on the generated track centerline, with the number of track segments set to 15. Table 13 presents the comparison of the geometric parameters between the generated and actual track centerlines. As shown in Table 13, the generated track’s length, height difference, and average slope are all close to those of the actual track, with maximum errors of 6%, 6%, and 6%, respectively. All slope values remain within the specified limits. As the weight of the height difference increases, the height difference of the generated centerline becomes closer to that of the actual track.
Figure 18 shows the track height variation along the distance for the generated centerlines. It can be observed that the height of the generated track centerline changes in accordance with the actual track trend. When the height difference weight is 0.7 or 0.9, the generated track height is close to the actual height. In contrast, when the weight is 0.5, the average slope of the generated track is smaller, resulting in lower track height beyond 600 m compared with those generated with height difference weights of 0.7 and 0.9.
Table 13: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.5$ 1625.8 123.7 No 0.0760
Weight factor $b=0.7$ 1666.5 115.4 No 0.0692
Weight factor $b=0.9$ 1634.3 117.7 No 0.0713
Actual track 1700 118 No 0.07
Figure 18: Track height variation along the distance for the generated centerlines.
5.2 Pyeongchange Track
Under the algorithm proposed in this study, an analysis was conducted on the proportionally scaled 2D data of the Pyeongchang track. The actual geometric parameters of the track are listed in Table 14 bi:14 .
Table 14: Actual geometric parameters of the Nagano track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1659 154 25 9.48
5.2.1 Track Generation
Under the proposed algorithm, a track generation analysis is conducted based on the uniformly scaled 2D data of the Nagano track. The calculation results are shown in Table 15. The generated track length, height difference, and average slope are all close to the actual values, with maximum errors of 1.0%, 1.0%, and 6.1%, respectively. Moreover, all slope values remain within the prescribed limits.
Figure 19 shows the track height variation along the distance for the generated centerlines. And the trends and values of different simulation results are similar and close to each other.
Table 15: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1647.8 152.0 No 0.0921
Generated track 2 1642.3 153.8 No 0.0935
Generated track 3 1641.6 154.0 No 0.0937
Actual track 1659 154 No 0.0981
Figure 19: Track height variation along the distance for the generated centerlines.
5.2.2 Influence of The Track Segment Count
For the Pyeongchang track, the influence of dividing the track into 11, 16, and 22 segments on the centerline generation results is analyzed.
Table 16 presents a comparison of the geometric parameters between the generated and actual track centerlines. As shown in Table 16, the track length, height difference, and average slope of the generated track under different numbers of track segments are all close to those of the actual track, with maximum errors of 1.0%, 4.8%, and 9.8%, respectively. Furthermore, all slope values remain within the specified limits.
Figure 20 shows the track height variation along the distance for the generated centerlines. Similar to the results of track generation in the actual 2D plane data section, Figure 20 demonstrates that the height of the generated centerline varies with distance in accordance with the actual track. A greater number of track segments results in more variations in local slopes while the overall change remains relatively small. This is because increasing the number of segments leads to more pronounced local slope fluctuations. Conversely, with fewer segments, the local variation is less significant, but the overall slope changes more substantially.
Table 16: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=11$ 1653.6 146.6 No 0.0885
Track segment count $s=16$ 1642.3 153.8 No 0.0935
Track segment count $s=22$ 1667.8 151.6 No 0.0907
Actual track 1659 154 No 0.0981
Figure 20: Track height variation along the distance for the generated centerlines.
5.2.3 Influence of The Weighting Factor
This section also analyzes the effect of the weight factor for height difference on the generated track centerline, with the number of track segments set to 16. Table 17 presents the comparison of the characteristic parameters between the generated and actual track centerlines. As shown in Table 17, the generated track’s length, height difference, and average slope are all close to those of the actual track, with maximum errors of 0.7%, 4.5%, and 10.0%, respectively. All slope values remain within the specified limits. As the weight of the height difference increases, the height difference of the generated centerline becomes closer to that of the actual track.
Figure 21 shows the track height variation along the distance for the generated centerlines. As shown in Figure 21, the height of the generated track centerline varies with track distance in a manner consistent with the trend of the actual track. When the track distance is between 0 and 600 m, the generated track heights under different height difference weights are close. However, after 400 m, due to the smaller average slope of the generated track, the track height generated with a height difference weight of 0.5 is lower than those generated with weights of 0.7 and 0.9.
Table 17: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.5$ 1662.4 147.0 No 0.0883
Weight factor $b=0.7$ 1647.8 152.0 No 0.0921
Weight factor $b=0.9$ 1650.0 155.3 No 0.0940
Actual track 1659 154.0 No 0.0981
Figure 21: Track height variation along the distance for the generated centerlines.
In summary, the proposed algorithm can effectively generate tracks for two cases with scaled 2D planar data. The generated track length, height difference, and average slope are close to the actual values, with maximum errors of 4.3%, 4.8%, and 8.6%, respectively. The slope values do not exceed the specified limits. On this basis, similar to the results of track generation with actual 2D plane data section, by changing the number of segments or adjusting the weight of height difference in the optimization equation, multiple stylistically different track centerlines can be generated. Within the parameter range used in this study, the maximum errors for track length, height difference, and average slope under different segmentation conditions are 4.4%, 4.8%, and 9.8%, respectively. Under different weight values for height difference, the maximum errors for track length, height difference, and average slope are 4.4%, 4.8%, and 10.0%, respectively.
6 Conclusions and future work
This study proposes a method for generating a 3D bobsleigh track centerline based on a 2D centerline. In accordance with international track design regulations, the method considers factors such as total track length, height, and slope continuity to establish a track generation optimization problem.
Within the selected range of tracks in this study, the proposed algorithm can compute various stylistic centerlines that are consistent with the trend of the actual track centerline, based on either actual or scaled 2D data. The generated centerlines are close to the actual values in terms of length, height difference, and average slope. The maximum errors calculated from actual and scaled 2D data are 1.7%, 3.2%, and 4.1% and 1.1%, 3.5%, and 4.3%, respectively. Moreover, all slope values remain within the specified limits.
By segmenting the track in different ways or adjusting the weight assigned to height difference in the optimization function, multiple stylistically distinct centerlines can be generated. The variety of track centerlines produced by the proposed algorithm can be reasonably utilized to assist in the design of bobsleigh track centerlines.
In the future, based on the wide range of generated national track data, the algorithm proposed in this study can be used to generate stylistically diverse bobsleigh tracks including track sections. Further work can focus on analyzing and constructing representative simulation tracks that incorporate key features from official tracks around the world. This will allow insights obtained from daily training, such as athletes’ driving strategies, to be more universally applicable across different official tracks worldwide. Moreover, exploring and comparing alternative optimization approaches including heuristic, stochastic, and learning-based methods may provide deeper insights into the trade-offs between solution quality, convergence speed, and robustness, relative to the current PGD-based framework.
7 Declaration of conflicting interests
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
8 Funding
The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work is supported by the Research on Open-Loop Dynamic Analysis Methods for Full-Vehicle Bobsleigh Systems of FAW R&D (Grant No. KH54633401).

### v2

@@
-PGD-based optimization of 3D bobsleigh track centerlines from 2D centerlines for simulation applications
+Projected gradient descent–based optimization of 3D bobsleigh track centerlines from 2D data for simulation
 Abstract
-The centerline of a bobsleigh track defines its geometry and is essential for simulation modeling.
-To reduce bBobsleigh training costs, leveraging the centerline of the bobsleigh track to construct a virtual environment that closely replicates real competitive settings presents a promising solution.
-However, publicly available centerline data are typically limited and it is imprecise to construct a training system solely based on 2-dimensional (2D) centerline.
-To address this practical issue, this paper proposes a method for generating a 3-dimensional (3D) track centerline based on 2D centerline data.
+This paper proposes a method for generating a 3-dimensional (3D) track centerline based on 2D centerline data.
 Incorporating international track design regulations, the method formulates an optimization problem that considers total track length, height difference, slope constraints, and geometric continuity.
 A Projected Gradient Descent (PGD) algorithm is used to solve the optimization problem.
-The generated 3D centerlines are compared with real track data, and the results show that the method can reproduce realistic centerline trends from original or scaled 2D data.
-For the selected track segment, the relative errors in total length, height difference, and average slope are within 1.7%, 3.2% and 4.1%, respectively, for real 2D data and within 1.1%, 3.5% and 4.3% respectively for scaled data.
-All slope values remain within the allowable limits.
-Moreover, by adjusting the segmentation or modifying the weight of height difference in the cost function, various centerline styles applicable to different competitions can be generated.
-Under different segmentation and weight factors, the maximum errors reach up to 4.4%, 4.8%, and 9.8%, and 4.4%, 4.8%, and 10.0%, respectively.
-The proposed method provides a flexible and efficient tool for supporting bobsleigh track centerline design.
+Within the selected trajectories and parameter ranges in this study, the proposed algorithm can compute centerlines whose characteristic parameter trends are consistent with those of the actual track, based on either actual or scaled 2D data.
+Compared with the actual track, the maximum errors ranges for length, height difference, and average gradient are 1.1%, 16.4%, and 13.5%.
+Track sensitivity to segmentation and height-difference weighting varies with data type, and appropriate selection of segmentation and weighting parameters effectively reduces deviations and enhances the geometric accuracy of the generated centerlines.
 keywords:
-Bobsleigh, track centerline generation, international Olympic bobsleigh competition rules, projected gradient descent
+Bobsleigh; track centerline generation; international Olympic bobsleigh competition rules; projected gradient descent
+articletype: ARTICLE TEMPLATE
 1 Introduction
-Since bobsleigh was formally recognized as a Winter Olympics event in 1924, it has attracted widespread attention from participating countries bi:bu4 .
-Bobsleighs are unpowered vehicles that rely entirely on the initial push from athletes and the conversion of gravitational potential energy into kinetic energy as they descend the track bi:1 ; 3 ; 5 .
-The competition is timed to an accuracy of milliseconds, and the final times often differ by only a few tenths of a second4 .
+Since bobsleigh was formally recognized as a Winter Olympics event in 1924, it has attracted widespread attention from participating countries (bi:bu4).
+Bobsleighs are unpowered vehicles that rely entirely on the initial push from athletes and the conversion of gravitational potential energy into kinetic energy as they descend the track (bi:1; 3; 5).
+The competition is timed to an accuracy of milliseconds, and the final times often differ by only a few tenths of a second(4).
 To minimize energy loss, athletes typically avoid using the brake unless necessary before crossing the finish line.
-The dynamics of the bobsleigh and the driver’s control skill are key factors influencing performance bi:bu1 ; r2 .
-However, due to restrictions from official governing bodies and the high cost of maintaining bobsleigh tracks7 , even countries with access to tracks often do not allow athletes to train on them regularly.
-Typically, limited test runs are permitted only shortly before official competitions bi:2 ; 1 ; bi:bu2 ; 6 .
+The dynamics of the bobsleigh and the driver’s control skill are key factors influencing performance (bi:bu1; roberts2013).
+However, due to restrictions from official governing bodies and the high cost of maintaining bobsleigh tracks(7), even countries with access to tracks often do not allow athletes to train on them regularly.
+Typically, limited test runs are permitted only shortly before official competitions (bi:2; rempfler2015; 4; 6).
 Consequently, simulation-based training and performance analysis have become vital tools for optimizing bobsleigh design and improving athlete technique.
 Researchers have developed virtual tracks, bobsleigh-blade/ice interaction models, multibody dynamics models of bobsleighs, and driver models.
-These simulation frameworks are increasingly being used to support bobsleigh structural design and help athletes devise optimal racing strategies.
-And there are limited studies on optimizing the centerline of bobsleigh tracks.
-Early bobsleigh simulation research was significantly constrained by the limited computational capabilities of CPUs at the time, often relying on simplified point-mass models moving on smooth hyperbolic surfaces bi:3 ; bi:4 .
-The study bi:4 developed a partial surface model of a bobsleigh track, treating the bobsleigh as a point mass to analyze its dynamic response under external forces.
+These simulation frameworks are increasingly being used to support bobsleigh structural design and help athletes devise optimal racing strategies (bi:2).
+Besides, there are limited studies on optimizing the centerlines of bobsleigh tracks.
+Early bobsleigh simulation research was significantly constrained by the limited computational capabilities of CPUs at the time, often relying on simplified point-mass models moving on smooth hyperbolic surfaces (bi:3; bi:4).
+The study (bi:4) developed a partial surface model of a bobsleigh track, treating the bobsleigh as a point mass to analyze its dynamic response under external forces.
 However, the authors also acknowledged that due to the lack of available data, virtual training environments differ significantly from the real settings, limiting the variety of technical strategies athletes can efficiently practice.
-Based on this point-mass framework and real data from the Lillehammer track in Norway, The study bi:5 further designed an optimal controller to simulate bobsleigh driving behavior, aiding in athlete training.
-To better capture the complex nonlinear dynamics of bobsleigh motion on narrow, twisted, banked, and icy tracks, a geometric model is proposed bi:6 , which represents the tracks as a sequence of discrete points arranged along a 3D spatial curve.
+Based on this point-mass framework and real data from the Lillehammer track in Norway, The study (bi:5) further designed an optimal controller to simulate bobsleigh driving behavior, aiding in athlete training.
+To better capture the complex nonlinear dynamics of bobsleigh motion on narrow, twisted, banked, and icy tracks, a geometric model is proposed (bi:6), which represents the tracks as a sequence of discrete points arranged along a 3D spatial curve.
 Moreover, a comprehensive simulation framework incorporating a blade–ice contact friction model is designed and validated on the Cesana and Whistler tracks, offering valuable guidance for track design and athlete training.
-The study bi:8 further validates its agreement between the simulated and experimental results on the Cesana track.
+The study (bi:8) further validates its agreement between the simulated and experimental results on the Cesana track.
 However, the study did not involve real-time simulation, and the applicability of the proposed model to other official tracks worldwide was not examined.
-Subsequently, the geometric model is further extended into a multibody dynamics model in bi:2 , where more detailed dynamic characteristics are considered, resulting in better driving fidelity.
-Recent advances in bobsleigh simulation have focused on improving blade–ice friction modeling and virtual training environments bi:bu3 ; bi:7 .
-The study bi:10 developed longitudinal and lateral friction models through experiments, FEM and nonlinear regression, validating their effectiveness in simulators to evaluate driving styles.
-Some studies further integrate steering and visual feedback or design immersive systems for training and tourism bi:3 ; bi:2 ; bi:10 ; bi:11 .
-The study bi:11 introduced a VR hammock system to mimic sliding sensations, while the studies bi:12 ; bi:13 proposed automated methods for generating 3D track surfaces.
+Subsequently, the geometric model is further extended into a multibody dynamics model (bi:2), where more detailed dynamic characteristics are considered, resulting in better driving fidelity.
+Recent advances in bobsleigh simulation have focused on improving blade–ice friction modeling and virtual training environments (bi:bu3; bi:7).
+The study (bi:10) developed longitudinal and lateral friction models through experiments, FEM and nonlinear regression, validating their effectiveness in simulators to evaluate driving styles.
+Some studies further integrate steering and visual feedback or design immersive systems for training and tourism (bi:2; bi:3; bi:10; bi:11).
+The study (bi:11) introduced a VR hammock system to mimic sliding sensations, while the studies (bi:12; bi:13) proposed automated methods for generating 3D track surfaces.
 However, most of the aforementioned studies focus on the simulation of only one or two tracks, and the conclusions drawn lack generalizability.
@@
 Therefore, acquiring a greater variety of available official track data worldwide has become an urgent need for enhancing both the diversity of research findings in simulation analysis and the general applicability of virtual training outcomes for bobsleigh athletes.
-The centerline of a bobsleigh track serves as a crucial dataset that defines the trajectory of the track.
-In contrast, the cross-sectional shapes are relatively standardized and can be generated using industrial software based on common templates bi:12 .
+The centerline of a bobsleigh track serves as a crucial dataset that defines a trajectory of a track.
+In contrast, the cross-sectional shapes are relatively standardized and can be generated using industrial software based on common templates (bi:12).
 This study focuses on typical bobsleigh track centerlines that determine the trajectory characteristics of the track and proposes a method for generating 3D centerlines from 2D centerlines.
 In accordance with international track design regulations, an optimization problem is formulated that considers factors such as total track length, height difference, slope, and continuity.
-A PGD is used to solve the problem.
+And a PGD algorithm is used to solve the problem.
 The generated 3D centerline is quantitatively compared with actual track data, and the influence of different track characteristics and algorithmic parameters on the generation results is investigated.
@@
 2.1 Formulation of the optimization problem
-According to the International Olympic Bobsleigh Competition Rules bi:14 , a bobsleigh track consists of a sequence of straight and curved sections.
+According to the International Olympic Bobsleigh Competition Rules (bi:14), a bobsleigh track consists of a sequence of straight and curved sections.
 These two types of track segments exhibit distinct geometric characteristics, which necessitate different approaches for their analysis and generation.
@@
 For a curved segment, when the distance $d$ between two points is extremely small, the segment between them can be approximated as a straight line.
-Therefore, if the coordinates of the previous point are $(x_{k-1},y_{k-1},z_{k-1})$, and the current point’s $x$ and $y$ coordinates are $(x_{k},y_{k})$, then the $z$-coordinate of the current point can be approximated as $z_{k}=z_{k}-d\cdot g_{k})$.
+Therefore, if the coordinates of the previous point are $(x_{k-1},y_{k-1},z_{k-1})$, and the current point’s $x$ and $y$ coordinates are $(x_{k},y_{k})$, then the $z$-coordinate of the current point can be approximated as $z_{k}=z_{k-1}-d\cdot g_{k}$.
 The distance $d$ is calculated as follows:
@@
 The generation of the 3D track from the 2D centerline can then be formulated as an optimization problem, as shown in Equation (3), in which the slope of each segment is optimized such that the resulting 3D coordinates satisfy the five aforementioned constraints.
-$J=\min_{g_{k}}\Bigg{[}$ $a\left(L-\sum_{k=2}^{n}d_{k}\right)+b\left(H-\sum_{k=2}^{n}\sqrt{(z_{k}-z_{k-1})^{2}}\right)$ (3)
-$+\frac{c}{n}\sum_{k=1}^{n}(K_{pk}-K_{sk})+d\left(\frac{1}{j}\sum_{m=1}^{j}g_{m}-\bar{g}\right)\Bigg{]}$
-[…]
+$J=\min_{g_{k}}\Bigg[$ $a\left(L-\sum_{k=2}^{n}d_{k}\right)^{2}+b\left(H-\sum_{k=2}^{n}\sqrt{(z_{k}-z_{k-1})^{2}}\right)^{2}$ (3)
+$+c\left(\frac{1}{n}\sum_{k=1}^{n}(K_{pk}-K_{sk})\right)^{2}+d\left(\frac{1}{j}\sum_{m=1}^{j}g_{m}-\bar{g}\right)^{2}\Bigg]$
+$\text{s.t.}\quad g_{k_{\text{min}}}\leq g_{k}\leq g_{k_{\text{max}}}$
 In this equation, $J$ represents the cost function to be minimized, and $j$ is the number of track segments.
-The parameters $a$, $b$, $c$, and $d$ are the weighting factors corresponding to the total length difference, height difference, curvature difference between the 2D and 3D centerlines, and the average slope difference, respectively.
-$g_{k_{\text{max}}}$ and $g_{k_{\text{min}}}$ denote the maximum and minimum allowable slopes.
-$d_{k}$ is the 3D distance between point $k$ and point $k-1$, calculated as follows:
+The parameters $a$, $b$, $c$, and $d$ are the weighting factors corresponding to the total length difference, height difference, curvature difference between the 2D and 3D centerlines, and the average slope difference, respectively, $g_{k_{\text{max}}}$ and $g_{k_{\text{min}}}$ denote the maximum and minimum allowable slopes, $d_{k}$ is the 3D distance between point $k$ and point k1k-1, calculated as follows:
 $d_{k}=\sqrt{(x_{k}-x_{k-1})^{2}+(y_{k}-y_{k-1})^{2}+(z_{k}-z_{k-1})^{2}}$ (4)
@@
 where $u=[\dot{x},\dot{y},\dot{z}]$ is the first derivative vector of the coordinates, and $v=[\ddot{x},\ddot{y},\ddot{z}]$ is the second derivative vector of the coordinates.
-To solve the bobsleigh track optimization problem established, a PGD bi:15 is used for parameter optimization.
+To solve the bobsleigh track optimization problem established, a PGD algorithm (9; 10) is used for parameter optimization.
 Compared to more complex methods such as genetic algorithms or simulated annealing, gradient descent offers faster convergence and requires fewer hyperparameters, making it more practical for repeated evaluations during track generation.
@@
 This approach does not rely on the explicit calculation of partial derivatives but instead approximates the gradient through a numerical difference method.
-More specifically, the direction of the gradient is approximated by evaluating the change in the cost function value between two successive iterations, rather than explicitly computing all partial derivatives.
-The underlying concept is similar to the finite difference method; however, it places greater emphasis on variations in the cost function itself rather than on exact gradients.
 The corresponding computation is given as follows:
@@
 $h_{i+1}=h_{i}-\eta\frac{\partial J}{\partial h_{i}}$ (8)
-where $\eta$ is the learning rate, which is set to $10^{-5}$.
-If the scaling factor is considered, it should also be updated using gradient descent as follows:
+where $\eta$ is the learning rate, which is set to $10^{-5}$. If the scaling factor is considered, it should also be updated using gradient descent as follows:
 $f_{i+1}=f_{i}-\eta\frac{\partial J}{\partial f_{i}}$ (9)
-Compared to SGD(Stochastic Gradient Descent) bi:sgd1 ; bi:sgd2 ; bi:sgd3 ; bi:sgd4 , PGD can be regarded as projecting the height difference of each segment during optimization to ensure that the corresponding slope of each segment falls within the prescribed limits, which in turn satisfies the overall constraint conditions.
+Compared to SGD (Stochastic Gradient Descent) algorithms (11; 12; 13), PGD algorithms can be regarded as projecting the height difference of each segment during optimization to ensure that the corresponding slope of each segment falls within the prescribed limits, which in turn satisfies the overall constraint conditions.
 Although the projection is not explicitly performed on each variable (e.g., the height difference of each segment) at every gradient descent step, it is effectively realized by adjusting the slope of each segment to meet the constraints.
@@
 where $d_{m}$ represents the total length of a specific segment of the track, and it is calculated as follows:
-$d_{m}$ $=\sum_{k=p+p_{2}+\cdots+p_{m-1}}^{p+p_{2}+\cdots+p_{m}}d_{k}$
+$d_{m}=\sum_{k=p_{1}+p_{2}+\cdots+p_{m-1}}^{p_{1}+p_{2}+\cdots+p_{m}}d_{k}$ (11)
 where $p_{m}$ denotes the number of points within segment $m$ of the track.
@@
 If $g_{m}$ exceeds this range (e.g., smaller than min slope or larger than max slope), the corresponding height difference $h_{m}$ is adjusted accordingly:
-{$h_{m}&=g_{k_{\text{min}}}\times d_{m},\quad&&g_{m}\leq g_{k_{\text{min}}}\\ h_{m}&=g_{k_{\text{max}}}\times d_{m},\quad&&g_{m}\geq g_{k_{\text{max}}}\end{aligned}\right.$ (11)
+$\left\{\begin{aligned} h_{m}&=g_{k_{\text{min}}}\times d_{m},\quad&&g_{m}\leq g_{k_{\text{min}}}\\ h_{m}&=g_{k_{\text{max}}}\times d_{m},\quad&&g_{m}\geq g_{k_{\text{max}}}\end{aligned}\right$. (12)
 $\Delta J=|J_{i+1}-J_{i}|<\delta$ (13)
@@
 2.2 Numerical Procedure
-Based on the optimization equation formulated in accordance with the track design rules and the PGD, the computational procedure is illustrated in the following figures.
-First, the initial segmental height differences and the corresponding cost function value are calculated.
-Then, using a randomly assigned initial gradient, the segmental height differences are updated, and the $z$-coordinates of all points are recalculated based on these updated values.
-Subsequently, the PGD algorithm enters an iterative loop.
+Based on the optimization equation formulated in accordance with the track design rules and the PGD algorithm, the computational procedure is illustrated in figure 2.
+Firstly, the initial segmental height differences and the corresponding cost function value are calculated.
+Then, using a assigned initial gradient to update segmental height differences, and the $z$-coordinates of all points are recalculated based on these updated values.
+Subsequently, the program enters the PGD algorithm iterative loop.
 At each iteration, it first checks whether the slope of each track segment falls within the specified bounds.
 If any segment’s slope exceeds the upper or lower limits, the projection method is applied to map it back into the feasible range.
-Using the projected slope values, the segmental height differences are updated, followed by recalculating the $z$-coordinates of all points.
+Using the projected slope values, to update the segmental height differences, followed by recalculating the $z$-coordinates of all points.
 The cost function is then evaluated again, and the height differences are further updated.
@@
 At that point, the final $z$ for all points on the track are determined.
-If a scaling factor is included in the optimization, the $x$ and $y$ are scaled at the beginning of each iteration, and the scaling factor is updated at the end of each iteration.
-The optimization process is illustrated in Figure 2.
-Figure 2: Calculation flow diagram.
-3 Results Analysis: Model Verification
+As shown in the dashed box in the figure, if a scaling factor is included in the optimization, the $x$ and $y$ are scaled at the beginning of each iteration, and the scaling factor is updated at the end of each iteration.
+Figure 2: Numerical procedure.
+3 Model Verification
 3.1 Track with Real 2D Plane Data
-Since some studies include real 2D plane data for a limited number of tracks bi:6 , this data can be used for solving the optimization problem.
+Since some studies include real 2D plane data for a limited number of tracks (bi:6), this data can be used for solving the optimization problem.
 The generated track data are then compared with the actual track data from the studies.
-In this section, the impact of the track’s ascent sections on the results is not considered.
-Therefore, only the data from the track’s descent sections are used for track generation and comparison analysis.
-In addition, the generation of the track centerline is a non-convex optimization problem, which may have multiple local optima.
-As a result, the algorithm may converge to different solutions in different runs.
-While the overall trend of the generated centerlines remains consistent, local variations may occur between runs.
-This characteristic enables the generation of diverse track designs, providing multiple alternative options for practical use and enhancing the flexibility and applicability of track design.
-Therefore, three random results are generated each time for comparing with the actual track centerline.
-3.1.1 Whistler Track
-The actual design geometric parameters of the Whistler track are shown in Table 1.
-Table 1: Actual geometric parameters of the Whistler track
-Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
-1278 149.3 20.4 11.6
-For the Whistler track, Figures 3 and 4 show the comparison of the actual track centerline’s 2D curvature and the actual and generated track centerlines’ 3D curvatures as a function of track distance.
-From Figures 3 and 4, it can be observed that there is no significant difference between the 2D curvature of the actual track centerline and the 3D curvature of both the actual and generated tracks across the track distance.
-Only small differences are found near the extreme values.
-Therefore, the impact of this factor is not further analyzed in the results.
-Figure 3: Comparison of the planar curvature and spatial curvature of the actual Whistler track centerline with track distance.
-Figure 4: Comparison of the 2D curvature and 3D curvature of the actual bobsleigh track centerline with track distance.
-Table 2 shows the comparison of the generated and actual Whistler bobsleigh track centerline geometric parameters.
-From Table 2, it can be observed that the generated track’s total length, height difference, and average slope are all close to those of the actual track, with maximum errors of 1.7%, 3.2%, and 4.1%, respectively.
-Additionally, the slope values do not exceed the defined limits.
-Table 2: Comparison of geometric parameters of the generated and actual Whistler track centerlines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Generated track 1 1291.1 150.9 No 0.1208
-Generated track 2 1290.2 144.5 No 0.1148
-Generated track 3 1290.7 148.4 No 0.1177
-Actual track 1278 149.3 No 0.1160
-Figure 5 shows the comparison of the track height variation along the distance for the generated and actual Whistler track centerlines.
-From Figure 5, it can be seen that the three generated track centerlines closely follow the trend of the actual track centerline, with the track height continuously decreasing as the track distance increases.
-Specifically, for Generated Track 1, the height is lower than the actual track up to approximately 400 m, but higher beyond this point.
-This is because the average slope of each segment beyond 400 m is higher than that of the actual track, resulting in a larger height difference.
-For Generated Track 3, the height closely matches the actual track up to approximately 100 m, then falls below the actual track between 100 m and 300 m, and exceeds the actual track between 300 m and 900 m, before aligning with the actual track again beyond 900 m.
-This is because the randomly assigned slopes and height differences are close to those of the actual track.
-Generated Track 2 is slightly lower than the actual track between 0–200 m and beyond 1025 m, but closely matches the actual track between 200 m and 1025 m.
-This is due to the randomly assigned slopes and height differences being lower than those of the actual track.
-Figure 5: Comparison of the track height variation along the distance for the generated and actual Whistler track centerlines.
-Figure 6 shows the comparison of the track slope variation along the distance for the generated and actual Whistler track centerlines.
-From Figure 6, it can be seen that according to the study bi:6 , at track distances of approximately 100m, the model calculated in this paper results in a centrifugal force that exceeds the maximum allowable centrifugal force.
-Although the track slope does not exceed the design slope value of 0.204, the higher average slope, shorter total length, and larger height difference contribute to the centrifugal force exceeding the maximum allowable value.
-The different generated track slopes in this algorithm exhibit consistent trends along the track distance, with only minor variations in amplitude.
-All generated track slopes are within the design value range, demonstrating that the algorithm can provide a rich variety of track generation results, and that relevant parameters can be constrained to meet the track design requirements.
-Figure 6: Comparison of the track slope variation along the distance for the generated and actual Whistler track centerlines.
-3.1.2 Censana Track
-Similar to the results of Whistler track, the actual geometric parameters of the Cesana track are shown in Table 3.
-Table 3: Actual geometric parameters of the Cesana track
-Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
-1359.2 124.5 18.3 9.1
-Table 4 shows the comparison of geometric parameters of the generated and actual Cesana track centerlines.
-From Table 4, the simulated values of the generated track’s total length, height difference, and average slope are all close to those of the actual track, with maximum errors of 0.81%, 1.2%, and 2.6%, respectively.
-Moreover, all slope values remain within the prescribed limits.
-Table 4: The comparison of geometric parameters of the generated and actual Cesana track centerlines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Generated track 1 1360.3 124.0 No 0.0934
-Generated track 2 1359.0 123.1 No 0.0926
-Generated track 3 1359.0 123.0 No 0.0930
-Actual track 1359.2 124.5 Exceeds by approximately 10m 0.0910
-Figure 7 shows the comparison between the track height along the centerline of the generated and actual Cesana tracks as a function of distance.
-From Figure 7, the three centerlines generated by the proposed algorithm all follow a similar trend to that of the actual track, with the track height decreasing as the distance along the track increases.
-For the generated Track 1, the height is close to that of the actual track before approximately 600 m, higher than the actual track between approximately 600 m and 1000 m, and lower than the actual track beyond approximately 1000 m.
-This is because the randomly assigned slope values in the segment between approximately 600 m and 1000 m are on average greater than those of the actual track.
-The result for generated Track 2 is similar to that of Track 1, with an even better match to the actual track before approximately 600 m.
-For generated Track 3, the result is similar to that of Track 1 before approximately 600 m, closely aligned with the actual track between approximately 600 m and 1000 m, but shows noticeable deviation after approximately 1000 m.
-This is due to the relatively smaller slope values assigned between approximately 600 m and 1000 m and the slightly lower overall height compared to the actual track.
-Figure 7: Comparison of track height variation along the distance for the generated and actual Cesana track centerlines.
-Figure 8 shows the comparison of track gradient along the centerlines between the generated and actual Cesana track.
-As shown in Figure 8, although the maximum designed gradient value is 0.183 bi:6 , the actual value exceeds 0.2, which may lead to a centrifugal force exceeding the allowable limit of 5g during cornering 8 .
-According to the calculation model presented in the study bi:6 , the maximum centrifugal force is approximately 4.5g, which can be attributed to the moderate height difference and sufficiently long total track length.
-While a lower gradient is critical for safety, an overly small gradient may reduce the differentiation in athlete performance, resulting in smaller performance gaps.
-The gradients of the tracks generated by the proposed algorithm follow a similar trend along the track distance, with only local variations in magnitude.
-All gradient values remain within the design limits, demonstrating that the gradient profile of the generated centerlines meets the required specifications.
-Figure 8: Comparison of the track slope variation along the distance for the generated and actual Cesana track centerlines.
-In summary, the proposed optimization equation can generate track centerlines with diverse characteristics based on real 2D data, and the trend of the generated centerline’s characteristic parameters is consistent with that of the actual track.
-The generated centerlines effectively cover the spatial domain of the original track centerlines.
-Simulation results on two selected tracks show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks, with maximum deviations of 1.7%, 3.2%, and 4.1%, respectively.
-Moreover, the gradients remain within the specified design limits, thereby validating the effectiveness of the proposed track centerline generation algorithm.
-The generation results are primarily influenced by parameters such as the randomly assigned segment gradients, height difference, distance, average gradient, and maximum gradient.
+Due to the fact that the actual track height difference parameter that can be obtained can only be reflected in the descending section of the track, and the descending section of the track is the key factor affecting the competition results.
+Therefore, in this section, the impact of the track’s final ascent sections on the results is not considered.
+Table 1 present comparisons of the geometric parameters of the generated and actual bobsleigh track centerlines for the Whistler and Cesana tracks.
+As shown in table 1, for the Whistler track, the generated tracks demonstrate strong consistency with the actual tracks in terms of total length, height difference, and average slope, with maximum deviations of 0.7%, 0.4%, and 0.5%, respectively.
+For the Cesana track, the corresponding deviations are 0.2%, 0.6%, and 1.1%.
+Furthermore, the generated tracks exhibit smaller average differences between the 2D and 3D curvature compared with their actual counterparts, while the slope values remain within the defined design limits.
+Table 1: Comparison of geometric parameters of the generated and actual Whistler track.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+Generated Track (Whistler) 1286.7 148.7 3.97×10-4 0.152 0.1182
+Actual Track (Whistler) 1278 149.3 9.76×10-4 0.204 0.1176
+Generated Track (Cesana) 1356.7 123.7 2.52×10-4 0.147 0.0920
+Actual Track (Cesana) 1359.2 124.5 5.04×10-4 0.183 0.0910
+Figure 3 shows the comparison of the track height variation and slope along the distance for the generated and actual track centerlines.
+Among them, Figure 3(a) is the Whistler track, and Figure 3(b) is the Cesana track.
+(a) Whistler track.
+(b) Cesana track.
+Figure 3: The comparison of the track height variation and slope along the distance for the generated and actual track centerlines.
+For the generated Whistler track, the height is lower than that of the actual track from 0 to approximately 245m, but higher between 245m and 1140m.
+Beyond 1140m, the heights converge.
+This is mainly attributed to the influence of slope variations.
+From the overall slope variation curve, the slopes of the generated track remain within the reasonable design range.
+The slopes of the actual and generated tracks alternate, resulting in height differences.
+For example, between 0 and 205m, the slope of the generated track is smaller than that of the actual track, leading to a lower height in this section.
+In contrast, from 205m to 245m, the slope of the generated track exceeds that of the actual track, causing the height to gradually approach that of the actual track.
+For the generated Cesana track, the height is close to that of the actual track from 0 to approximately 115m, but higher between 115m and 1266m.
+Beyond 1266m, the heights converge as well.
+In addition, due to slope variations, the slope of the generated track is greater than that of the actual track in most regions between 115m and 950m, resulting in higher track heights.
+In contrast, between 950m and 1110m, the slope of the generated track is smaller than that of the actual track, leading to a gradual convergence of height toward that of the actual track.
+According to the study by bi:6, for the Whistler track, when the track distance is approximately 100m, the centrifugal force of the actual track exceeds the maximum allowable value.
+Although the track slope does not exceed the design limit of 0.204, a higher average slope, shorter total length, and larger height difference contribute to the centrifugal force surpassing the permissible threshold of 5g.
+In the case of the Cesana track, although its maximum design slope is 0.183, the actual slope exceeds 0.2, which may cause the centrifugal force during turns to exceed the allowable limit.
+The maximum centrifugal force of the actual track is about 4.5g, which remains within a reasonable range owing to the appropriate height difference and sufficient total track length.
+Although a smaller slope is critical for safety, an excessively small slope may reduce performance differences among athletes, thereby narrowing performance gaps.
+Figures 4 shows the comparison of trajectory differences between the generated and actual track centerlines, in terms of 2D and 3D curvature variation along the track distance.
+Among them, Figure 4(a) is the Whistler track, and Figure 4(b) is the Cesana track.
+As shown in the figures, in most regions, the differences between the 2D and 3D curvature of the generated track are smaller than or close to those of the actual track.
+This prevents excessive fluctuations in the slope in the z-direction and mitigates large vertical accelerations, thereby improving the safety and smoothness of the track design.
+(a) Whistler track.
+(b) Cesana track.
+Figure 4: The comparison of trajectory differences between the generated and actual track centerlines.
+In summary, the proposed optimization equation can generate track centerlines based on real 2D data, and the proposed algorithm can compute centerlines whose characteristic parameter trends are consistent with those of the actual track.
+Simulation results on two selected tracks show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks.
+Moreover, the slopes remain within the specified design limits, thereby validating the effectiveness of the proposed track centerline generation algorithm.
+The generation results are primarily influenced by parameters such as height difference, distance, average gradient, maximum gradient, average differences between the 2D and 3D curvature.
 In addition, the two real tracks analyzed herein highlight potential directions for further optimization.
@@
 Taking the Whistler track as an example in this section, the optimization problem considers a scaling factor $f$, where the horizontal $x$ and $y$ coordinates of the Whistler track are uniformly scaled to 90% of their original values (isotropic scaling is applied, with no inconsistency between the scaling ratios of $x$ and $y$ coordinates).
-Figure 9 presents a comparison between the generated and actual $x$ and yy coordinates of the track centerline.
+Figure 5 presents the comparison between the generated and actual $x$ and $y$ coordinates of the track centerlines.
 As shown in figure 5, the $x$ and $y$ coordinates calculated by the proposed algorithm follow the same trend as the actual values.
-The computed scaling factors $f$ are 1.1102, 1.1196, and 1.0857, corresponding to relative errors of 0.1%, 7.6%, and 2.3%, respectively (with respect to $f$ = 1.1).
-These results demonstrate that the proposed algorithm can effectively estimate the original $x$ and $y$ coordinates of the track centerline prior to scaling.
-Figure 9: Comparison of the track slope variation along the distance for the generated and actual Cesana track centerlines.
+The computed scaling factor $f$ is 1.1126, corresponding to a relative error of 0.14% (with respect to $f$ = 1.1111).
+The result demonstrates that the proposed algorithm can effectively estimate the original $x$ and $y$ coordinates of the track centerline prior to scaling.
+Figure 5: The comparison between the generated and actual $x$ and $y$ coordinates of the track centerlines.
 To further analyze the computational results, Table 2 presents a comparison between the geometric parameters of the generated and actual track centerlines.
-As shown in Table 5, the simulated values of track length, height difference, and average slope of the generated track are all close to those of the actual track, with maximum errors of 1.1%, 3.5%, and 4.3%, respectively.
-Additionally, all slope values remain within the prescribed limits.
-Table 5: The comparison of geometric parameters of the generated and actual track centerlines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Generated track 1 1291.2 146.9 No 0.1150
-Generated track 2 1290.8 143.9 No 0.1126
-Generated track 3 1292.2 154.6 No 0.1210
-Actual track 1278 149.3 No 0.1160
-Figure 10: Comparison of track height variation along the distance for the generated and actual track centerlines.
-Figure 11: Comparison of the track slope variation along the distance for the generated and actual track centerlines.
-Figures 10 and 11 respectively compare the variation of track height and slope along the distance between the generated and actual track centerlines.
-Similar to the results in model verification section, the height variation trends of the centerlines generated by the proposed algorithm are consistent with those of the actual track, and all slope values of the generated centerlines remain within the design limits.
-In summary, the proposed algorithm can compute the original (pre-scaled) $x$ and $y$ coordinates from the scaled $x–y$ plane coordinates.
-Similar to the algorithm based on the original unscaled $x–y$ plane coordinates, the one that incorporates the scaling factor is also capable of generating track centerlines with diverse characteristics while maintaining consistency with the overall trends of the actual track centerline.
-The generated centerlines sufficiently cover the range of the original centerline.
-Validation using the selected track demonstrates that the generated track’s length, height difference, and average slope are all close to the actual values, with maximum errors of 1.1%, 3.5%, and 4.3%, respectively.
-Moreover, the slope values do not exceed the specified limits, thereby verifying the effectiveness of the proposed track generation algorithm.
-The track generation results are mainly influenced by parameters such as the randomly assigned slope of each segment, the track’s height difference, length, average slope, and maximum slope.
-4 Results Analysis: Influence of Different Parameters on Track Generation with Actual 2D Plane Data
-In this section, the Whistler and Cesana tracks are selected as case studies to analyze the effects of segment count and weighting factors on track generation under the condition of actual 2D plane data.
-4.1 Whistler Track
-4.1.1 Influence of Track Segment Count
-For the Whistler track, the effect of dividing the track into 15, 26, and 33 segments on the generation results was analyzed.
-Table 6 presents a comparison of the geometric parameters between the generated and actual Whistler track centerlines.
-As shown in Table 6, under different segmentation settings, the simulated values of track length, height difference, and average slope are all close to those of the actual track, with maximum errors of 1.1%, 4.8%, and 7.7%, respectively.
-Additionally, all slope values remain within the specified limits.
-Figure 12 shows the comparison of height variation along the distance between the generated and actual track centerlines.
-It can be observed that the height profiles of the generated centerlines follow the same trend as that of the actual track.
-For the Whistler track, increasing the number of track segments results in greater local slope variation, but the overall variation remains relatively minor.
-This is due to the increased number of local slope changes caused by a higher segmentation count.
-Conversely, when the number of segments is lower, local variations are reduced, while the overall slope variation becomes more pronounced, leading to better agreement with the actual track.
-In this case, as the track distance increases, the height also increases.
-In the 0–600 m segment of the track, the generated centerlines based on 15 and 33 segments show higher consistency with the actual track.
-In the 600–1200 m segment, the track centerline generated with 15 segments aligns more closely with the actual track.
-This is because in the 0–600 m segment, the actual centerline exhibits a relatively steep slope, so a larger slope value in the generated centerline leads to a closer match with the actual track.
-In contrast, in the 600–1200 m segment, the actual centerline has a gentler slope, so a smaller slope value in the generated centerline results in better agreement with the actual track.
-Table 6: Comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Track segment count $s=15$ 1291.5 151.4 No 0.1198
-Track segment count $s=26$ 1291.7 156.5 No 0.1249
-Track segment count $s=33$ 1291.7 156.2 No 0.1222
-Actual track 1278 149.3 No 0.1160
-Figure 12: Comparison of track height variation along the distance for the generated and actual track centerlines.
-4.1.2 Influence of The Weighting Factor
-The height difference largely determines the elevation change of each segment along the track.
-Therefore, this section analyzes the influence of the weighting factor for height difference on the centerline generation results.
-The number of track segments is set to 26.
-Table 7 presents the comparison of geometric parameters between the generated and actual track centerlines.
-As shown in Table, the simulated values of track length, height difference, and average gradient under different height difference weighting factors are all close to the actual values.
-The maximum errors are 0.7%, 1.7%, and 3.8%, respectively.
-Furthermore, all slope values remain within the specified limits.
-As the height difference weighting increases, the height difference of the generated centerline becomes more consistent with the actual value.
-Figure 13 shows a comparison between the generated and actual centerline track height variation with distance.
-As illustrated in Figure 13, the variation trend of the generated centerline height is consistent with that of the actual track.
-For the segment between 200 m and 600 m, the generated centerlines with weighting factors of 0.7 and 0.9 demonstrate a higher degree of agreement with the actual data.
-For the segment between 600 m and 1000 m, those with weighting factors of 0.5 and 0.7 are more consistent with the actual centerline.
-This is because for similar track lengths, smaller height differences generally correspond to gentler gradients.
-In the segment between 200 m and 600 m, the generated centerline with a larger gradient shows better agreement with the actual track, whereas in the segment from 600 m to 1000 m, the centerline with a smaller gradient corresponds more closely to the actual values.
-Table 7: Comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Weight factor $b=0.5$ 1287.6 151.9 No 0.1183
-Weight factor $b=0.7$ 1287.6 151.4 No 0.1197
-Weight factor $b=0.9$ 1287.5 151.1 No 0.1204
-Actual track 1278 149.3 No 0.1160
-Figure 13: Comparison of track height variation along the distance for the generated and actual track centerlines.
-4.2 Cesana Track
-4.2.1 Influence of Track Segment Count
-For the Cesana track, the influence of dividing the track into 12, 21, and 29 segments on the centerline generation results is analyzed.
-Table 8 presents the comparison of geometric parameters between the generated and actual track centerlines.
-As shown in Table 8, under different numbers of segments, the simulated values of track length, height difference, and average gradient are all close to the actual values, with maximum errors of 0.2%, 0.9%, and 2.4%, respectively.
-In all cases, the gradient remains within the specified limits.
-Figure 14 shows the variation of track height along the distance for both the generated and actual centerlines.
-As can be seen, the generated centerline height follows a trend consistent with the actual track.
-For the segments from 0 to 600 m and after 1000 m, the centerline generated with 29 segments exhibits the highest degree of agreement with the actual centerline.
-In contrast, for the segment from 600 m to 1000 m, the centerline generated with 12 segments shows the highest level of agreement.
-This is because increasing the number of segments leads to more local variations in the track gradient, while the overall variation remains limited.
-Conversely, a smaller number of segments results in fewer local variations but more pronounced overall gradient changes.
-For this particular track, in the segments from 0 to 600 m and after 1000 m, where the actual centerline has relatively steep gradients, the generated centerline better matches the actual one when the gradient values are higher.
-In the segment from 600 m to 1000 m, where the actual track gradient is gentler, the generated centerline more closely matches the actual track when the gradient values are smaller.
-Table 8: Comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Track segment count $s=12$ 1356.6 123.4 No 0.0932
-Track segment count $s=21$ 1356.4 124.0 No 0.0932
-Track segment count $s=29$ 1356.6 123.9 No 0.0931
-Actual track 1359.2 124.5 No 0.0910
-Figure 14: Comparison of track height variation along the distance for the generated and actual track centerlines.
-4.2.2 Influence of Weighting Factor
-This section also analyzes the influence of the height difference weight factor on the results of centerline generation, with the number of track segments set to 12.
-Table 9 presents the comparison of geometric parameters between the generated and actual track centerlines.
-As shown in Table 9, under different weight factors for height difference, the generated track length, height difference, and average slope are all close to those of the actual track, with maximum errors of 8%, 9%, and 5%, respectively, and all slope values remain within the allowable range.
-Figure 15 shows the comparison of track height versus distance between the generated and actual track centerlines.
-As seen in Figure 15, in the section of 0–600 m, the generated track centerline with a weight factor of 0.1 exhibits better agreement with the actual centerline.
-In the section of 600–1300 m, the generated centerlines with weight factors of 0.7 and 0.9 show better agreement with the actual track.
-This is because for similar track distances, a smaller height difference generally results in a smoother profile as indicated by the average slope.
-In the 0–600 m section, a larger slope in the generated track centerline better matches the actual track, while in the 600–1300 m section, a smaller slope in the generated centerline aligns more closely with the actual track centerline.
-Table 9: Comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Weight factor $b=0.1$ 1357.1 128.9 No 0.0974
-Weight factor $b=0.7$ 1356.8 125.0 No 0.0940
-Weight factor $b=1.3$ 1356.7 124.7 No 0.0941
-Actual track 1359.2 124.5 No 0.0910
-Figure 15: Comparison of track height variation along the distance for the generated and actual track centerlines.
-To summarize, by dividing the track into different numbers of segments or adjusting the weight of height difference in the optimization equation, multiple centerlines with distinct characteristics can be generated.
-Within the parameter ranges used in this study, the generated track length, height difference, and average slope under different segmentations are all close to those of the actual tracks, with maximum errors of 1.1%, 4.8%, and 7.7%, respectively.
-Under different height difference weights, the maximum errors in track length, height difference, and average slope are 0.7%, 3.4%, and 7.0%, respectively.
-Moreover, all generated slopes remain within the prescribed limits.
-The agreement between the generated and actual centerlines varies depending on the track and the segmentation or weight factor used.
-The variety of centerlines generated by the proposed algorithm can be flexibly applied to meet specific track design requirements, thereby serving as a valuable aid in centerline design.
-5 Results Analysis: Influence of Different Parameters on Track Generation under Scaled 2D Planar Data
-In this section, the Nagano track in Japan and the Pyeongchang track in South Korea are selected as case studies to analyze the influence of segmentation and weight factors on track generation based on scaled 2D planar data.
-5.1 Nagano Track
-This section selects the Nagano track and the PyeongChang track as case studies to analyze the influence of the number of segments and the weight coefficient on the generated track under conditions involving scaled 2D planar data.
-Since the starting point of the final ascending segment of the track cannot be determined and the slope design rules for the ascending segment are similar to those for the descending segment bi:14 , the final ascending segment is treated as part of the descending segment.
-Its height change is included in the total height difference for analysis.
-Under the algorithm proposed in this study, an analysis was conducted on the proportionally scaled 2D data of the Nagano track.
-The actual geometric parameters of the track are listed in Tables 10 bi:14 .
-Table 10: Actual geometric parameters of the Nagano track
-Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
-1700 118 15 7
-5.1.1 Track Generation
-Under the proposed algorithm, a track generation analysis is conducted based on the uniformly scaled 2D data of the Nagano track in Japan.
-The calculation results are shown in Table 11.
-Similar to the results in model verification section, the generated track length, height difference, and average slope are all close to the actual values, with maximum errors of 4.3%, 4.8%, and 8.6%, respectively.
-Moreover, all slope values remain within the prescribed limits.
-Table 11: Comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Generated track 1 1625.8 123.7 No 0.0760
-Generated track 2 1652.4 115.6 No 0.0699
-Generated track 3 1658 113.9 No 0.0686
-Actual track 1700 118 No 0.0700
-Figure 16: Comparison of track height variation along the distance for the generated and actual track centerlines.
-5.1.2 Influence of Track Segment Count
-For the Nagano track, the influence of dividing the track into 10, 15, and 21 segments on the centerline generation results is analyzed.
-Table 12 presents a comparison of the geometric parameters between the generated and actual track centerlines.
-As shown Table 12, the track length, height difference, and average slope of the generated track under different numbers of track segments are all close to those of the actual track, with maximum errors of 4.4%, 3.6%, and 8.6%, respectively.
-Furthermore, all slope values remain within the specified limits.
-Figure 18 shows the track height variation along the distance for the generated centerlines.
-Similar to the results of track generation with actual 2D plane data section, Figure 18 demonstrates that the height of the generated centerline varies with distance in accordance with the actual track.
-A greater number of track segments results in more variations in local slopes while the overall change remains relatively small.
-This is because increasing the number of segments leads to more pronounced local slope fluctuations.
-Conversely, with fewer segments, the local variation is less significant, but the overall slope changes more substantially.
-Table 12: Comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Track segment count $s=10$ 1653.3 113.8 No 0.0688
-Track segment count $s=15$ 1625.8 123.7 No 0.076
-Track segment count $s=21$ 1656.4 117.3 No 0.0708
-Actual track 1700 118 No 0.07
-Figure 17: Track height variation along the distance for the generated centerlines.
-This section also analyzes the effect of the weight factor for height difference on the generated track centerline, with the number of track segments set to 15.
-Table 13 presents the comparison of the geometric parameters between the generated and actual track centerlines.
-As shown in Table 13, the generated track’s length, height difference, and average slope are all close to those of the actual track, with maximum errors of 6%, 6%, and 6%, respectively.
-All slope values remain within the specified limits.
-As the weight of the height difference increases, the height difference of the generated centerline becomes closer to that of the actual track.
-Figure 18 shows the track height variation along the distance for the generated centerlines.
-It can be observed that the height of the generated track centerline changes in accordance with the actual track trend.
-When the height difference weight is 0.7 or 0.9, the generated track height is close to the actual height.
-In contrast, when the weight is 0.5, the average slope of the generated track is smaller, resulting in lower track height beyond 600 m compared with those generated with height difference weights of 0.7 and 0.9.
-Table 13: The comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Weight factor $b=0.5$ 1625.8 123.7 No 0.0760
-Weight factor $b=0.7$ 1666.5 115.4 No 0.0692
-Weight factor $b=0.9$ 1634.3 117.7 No 0.0713
-Actual track 1700 118 No 0.07
-Figure 18: Track height variation along the distance for the generated centerlines.
-5.2 Pyeongchange Track
-Under the algorithm proposed in this study, an analysis was conducted on the proportionally scaled 2D data of the Pyeongchang track.
-The actual geometric parameters of the track are listed in Table 14 bi:14 .
-Table 14: Actual geometric parameters of the Nagano track
-Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
-1659 154 25 9.48
-5.2.1 Track Generation
-Under the proposed algorithm, a track generation analysis is conducted based on the uniformly scaled 2D data of the Nagano track.
-The calculation results are shown in Table 15.
-The generated track length, height difference, and average slope are all close to the actual values, with maximum errors of 1.0%, 1.0%, and 6.1%, respectively.
-Moreover, all slope values remain within the prescribed limits.
-Figure 19 shows the track height variation along the distance for the generated centerlines. And the trends and values of different simulation results are similar and close to each other.
-Table 15: The comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Generated track 1 1647.8 152.0 No 0.0921
-Generated track 2 1642.3 153.8 No 0.0935
-Generated track 3 1641.6 154.0 No 0.0937
-Actual track 1659 154 No 0.0981
-Figure 19: Track height variation along the distance for the generated centerlines.
-5.2.2 Influence of The Track Segment Count
-For the Pyeongchang track, the influence of dividing the track into 11, 16, and 22 segments on the centerline generation results is analyzed.
-Table 16 presents a comparison of the geometric parameters between the generated and actual track centerlines.
-As shown in Table 16, the track length, height difference, and average slope of the generated track under different numbers of track segments are all close to those of the actual track, with maximum errors of 1.0%, 4.8%, and 9.8%, respectively.
-Furthermore, all slope values remain within the specified limits.
-Figure 20 shows the track height variation along the distance for the generated centerlines.
-Similar to the results of track generation in the actual 2D plane data section, Figure 20 demonstrates that the height of the generated centerline varies with distance in accordance with the actual track.
-A greater number of track segments results in more variations in local slopes while the overall change remains relatively small.
-This is because increasing the number of segments leads to more pronounced local slope fluctuations.
-Conversely, with fewer segments, the local variation is less significant, but the overall slope changes more substantially.
-Table 16: The comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Track segment count $s=11$ 1653.6 146.6 No 0.0885
-Track segment count $s=16$ 1642.3 153.8 No 0.0935
-Track segment count $s=22$ 1667.8 151.6 No 0.0907
-Actual track 1659 154 No 0.0981
-Figure 20: Track height variation along the distance for the generated centerlines.
-5.2.3 Influence of The Weighting Factor
-This section also analyzes the effect of the weight factor for height difference on the generated track centerline, with the number of track segments set to 16.
-Table 17 presents the comparison of the characteristic parameters between the generated and actual track centerlines.
-As shown in Table 17, the generated track’s length, height difference, and average slope are all close to those of the actual track, with maximum errors of 0.7%, 4.5%, and 10.0%, respectively.
-All slope values remain within the specified limits.
-As the weight of the height difference increases, the height difference of the generated centerline becomes closer to that of the actual track.
-Figure 21 shows the track height variation along the distance for the generated centerlines.
-As shown in Figure 21, the height of the generated track centerline varies with track distance in a manner consistent with the trend of the actual track.
-When the track distance is between 0 and 600 m, the generated track heights under different height difference weights are close.
-However, after 400 m, due to the smaller average slope of the generated track, the track height generated with a height difference weight of 0.5 is lower than those generated with weights of 0.7 and 0.9.
-Table 17: The comparison of geometric parameters of the generated and actual track centrelines
-Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
-Weight factor $b=0.5$ 1662.4 147.0 No 0.0883
-Weight factor $b=0.7$ 1647.8 152.0 No 0.0921
-Weight factor $b=0.9$ 1650.0 155.3 No 0.0940
-Actual track 1659 154.0 No 0.0981
-Figure 21: Track height variation along the distance for the generated centerlines.
-In summary, the proposed algorithm can effectively generate tracks for two cases with scaled 2D planar data.
-The generated track length, height difference, and average slope are close to the actual values, with maximum errors of 4.3%, 4.8%, and 8.6%, respectively.
-The slope values do not exceed the specified limits.
-On this basis, similar to the results of track generation with actual 2D plane data section, by changing the number of segments or adjusting the weight of height difference in the optimization equation, multiple stylistically different track centerlines can be generated.
-Within the parameter range used in this study, the maximum errors for track length, height difference, and average slope under different segmentation conditions are 4.4%, 4.8%, and 9.8%, respectively.
-Under different weight values for height difference, the maximum errors for track length, height difference, and average slope are 4.4%, 4.8%, and 10.0%, respectively.
+As shown in Table 2, Simulation results show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks, with maximum deviations of 0.87%, 3.8%, and 3.6%, respectively.
+Similar to Section 3.1, the generated tracks exhibit smaller average differences between the 2D and 3D curvature compared with their actual counterparts, while the slope values remain within the defined design limits.
+Table 2: Comparison of geometric parameters of the generated and actual Whistler track.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+Generated Track (Whistler) 1289.1 154.9 4.17×10-4 0.1710 0.1218
+Actual Track (Whistler) 1278 149.3 9.76×10-4 0.204 0.1176
+In summary, the results obtained after applying the scaling factor further demonstrate the robustness and generalizability of the proposed optimization equation.
+Although minor deviations arise due to the scaling process, the overall geometric characteristics of the generated centerlines remain consistent with those of the actual tracks.
+The trends of the characteristic parameters—such as height difference, total length, and average gradient—are preserved, and the slopes stay within the defined design limits.
+These findings indicate that the proposed track centerline generation algorithm maintains accuracy and stability under scaled conditions, thereby confirming its applicability to tracks of varying dimensions.
+Moreover, the analysis of the scaled cases provides valuable insights for improving the parameter selection and enhancing the adaptability of the algorithm.
+4 Influence of Different Parameters on Track Generation
+In this section, tracks with real 2D centerline data (Whistler and Cesana tracks) and scaled 2D data (Nagano and Pyeongchang tracks) are selected as case studies to analyze the effects of segment count and weighting factors on track generation under the condition of actual and scaled 2D plane data.
+And the number of track segments and learning rate are determined separately for each track.
+4.1 Influence of the Number of Track Segments
+Figure 6 is influence of the number of track segments on track height variation along the track distance.
+As shown in figure 6, the proposed method reproduces the global geometric characteristics of each track with small deviations, demonstrating stable optimization performance under different segmentation conditions.
+Simulation results for all four tracks indicate that the maximum deviation ranges for length, height difference, and average gradient are 0.1%-1.1%, 1.3%-13.6% and 1.1%-10.7%, respectively, with detailed values provided in Appendix A (Table 36).
+All slope values remain within the specified design limits, confirming the robustness of the optimization framework under different weighting conditions.
+Increasing the number of segments primarily provides greater local flexibility, allowing more detailed adjustments of the segment-wise gradients.
+However, the effect of segmentation varies among tracks.
+For tracks with real 2D centerline data, show greater sensitivity to segmentation, as finer divisions better capture their local slope changes.
+In contrast, for tracks using scaled 2D data are less affected by segmentation count, showing stable results across different settings.
+Overall, these results suggest that the sensitivity to segmentation differs across tracks, and that selecting an appropriate number of track segments can effectively reduce deviations between the generated and actual track data, improving the geometric accuracy of the reconstructed centerlines.
+Figure 6: Influence of the number of track segments on track height variation along the track distance.
+4.2 Influence of the Weighting Factor
+The weighting factor of the height-difference term determines the degree to which the optimization prioritizes vertical conformity between the generated and actual tracks.
+In this section, its influence is analyzed across four representative tracks under varying weighting values (0.5, 1.0, and 1.5).
+Detailed values are provided in Appendix A (Table 710)
+Figure 7 is influence of the weighting factor on track height variation along the track distance.
+As shown in figure 7, Across all cases, the proposed algorithm produces track centerlines whose length, height difference, and average gradient closely match those of the actual tracks, with maximum deviation ranges of 0.2–0.7%, 1.0–16.4%, and 0.8–13.5%, respectively.
+Similar to the results obtained for the number of track segments, all slope values remain within the specified design limits.
+However, the sensitivity to the height-difference weighting varies among tracks.
+For tracks with real 2D centerline data, increasing the weighting factor enhances the alignment of the generated elevation profile with the actual track, as the optimization places greater emphasis on matching the total height difference.
+In contrast, for tracks using scaled 2D data, the effect of the weighting factor is less pronounced, since their geometric variation is already constrained by the scaling ratio.
+Consequently, the influence of the height-difference weighting is track-dependent, and selecting an appropriate weighting factor can effectively reduce deviations between the generated and actual track data, further improving the overall accuracy of the reconstructed 3D centerlines.
+Figure 7: Influence of the weighting factor on track height variation along the track distance.
 5 Conclusions and future work
@@
 In accordance with international track design regulations, the method considers factors such as total track length, height, and slope continuity to establish a track generation optimization problem.
-Within the selected range of tracks in this study, the proposed algorithm can compute various stylistic centerlines that are consistent with the trend of the actual track centerline, based on either actual or scaled 2D data.
-The generated centerlines are close to the actual values in terms of length, height difference, and average slope.
-The maximum errors calculated from actual and scaled 2D data are 1.7%, 3.2%, and 4.1% and 1.1%, 3.5%, and 4.3%, respectively.
-Moreover, all slope values remain within the specified limits.
-By segmenting the track in different ways or adjusting the weight assigned to height difference in the optimization function, multiple stylistically distinct centerlines can be generated.
-The variety of track centerlines produced by the proposed algorithm can be reasonably utilized to assist in the design of bobsleigh track centerlines.
+Within the selected range of tracks in this study, the proposed algorithm can compute centerlines whose characteristic parameter trends are consistent with those of the actual track, based on either actual or scaled 2D data.
+Compared with the actual track, the generated centerlines are close to the actual values in terms of length, height difference, and average slope.
+Simulation results for all four tracks indicate that the maximum deviation ranges for length, height difference, and average gradient are 0.1%-1.1%, 1.0%-16.4%, and 0.8%-13.5%, respectively.
+The sensitivity to segmentation and height-difference weighting differs across tracks, for tracks with real 2D centerline data, higher sensitivity to segmentation and weighting factors is observed, as finer divisions better capture local slope variations, and increasing the weighting factor improves the alignment between the generated and actual elevation profiles.
+In contrast, tracks based on scaled 2D data exhibit lower sensitivity to both segmentation count and weighting factor, maintaining consistent results under different configurations.
+Selecting an appropriate number of track segments and weighting factors can effectively reduce deviations between the generated and actual track data, improving the geometric accuracy of the reconstructed centerlines.
 In the future, based on the wide range of generated national track data, the algorithm proposed in this study can be used to generate stylistically diverse bobsleigh tracks including track sections.
@@
 This will allow insights obtained from daily training, such as athletes’ driving strategies, to be more universally applicable across different official tracks worldwide.
-Moreover, exploring and comparing alternative optimization approaches including heuristic, stochastic, and learning-based methods may provide deeper insights into the trade-offs between solution quality, convergence speed, and robustness, relative to the current PGD-based framework.
+Moreover, exploring and comparing alternative optimization approaches including heuristic, stochastic, and learning-based methods may provide deeper insights into the trade-offs between solution quality, convergence speed, and robustness, relative to the current algorithm based framework.
 6 Declaration of conflicting interests
@@
 KH54633401).
+Data availability
+The data that support the findings of this study are available from the corresponding author, Jicheng Chen, upon reasonable request.
+Appendix A Calculation results
+Table 3: Comparison of geometric parameters of the generated and actual Whistler track under different numbers of track segments.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+s=15 1286.8 147.3 $4.52\times 10^{-4}$ 0.1703 0.1198
+s=26 1286.7 148.2 $3.79\times 10^{-4}$ 0.156 0.1187
+s=33 1286.7 148.7 $3.97\times 10^{-4}$ 0.152 0.1182
+Actual Track (Whistler) 1278 149.3 $9.76\times 10^{-4}$ 0.204 0.1176
+Table 4: Comparison of geometric parameters of the generated and actual Cesana track under different numbers of track segments.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+s=10 1356.5 122.2 1.94×10-4 0.1425 0.0939
+s=15 1356.4 123.3 1.91×10-4 0.125 0.0926
+s=21 1356.7 123.7 2.52×10-4 0.147 0.0920
+Actual Track (Cesana) 1359.2 124.5 5.04×10-4 0.183 0.0910
+Table 5: Comparison of geometric parameters of the generated and actual Nagano track under different numbers of track segments.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+s=10 1702.0 124.5 2.70×10-4 0.1057 0.0734
+s=12 1701.5 124.6 1.95×10-4 0.0867 0.0734
+s=15 1699.7 126.5 2.04×10-4 0.0878 0.0746
+Actual Track (Nagano) 1700 118 - 0.150 0.0700
+Table 6: Comparison of geometric parameters of the generated and actual Pyeongchang track under different numbers of track segments.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+s=11 1662.0 166.5 2.47×10-4 0.1315 0.1008
+s=16 1648.6 170.3 1.87×10-4 0.1170 0.1038
+s=22 1677.3 175.0 1.71×10-4 0.1150 0.1049
+Actual Track (Pyeongchang) 1659 154 - 0.25 0.0948
+Table 7: Comparison of geometric parameters of the generated and actual Whistler track under different height-difference weighting factors.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+b=0.5 1287.0 150.8 4.02×10-4 0.1434 0.1167
+b=1 1286.7 148.7 3.97×10-4 0.152 0.1182
+b=1.5 1286.7 149.1 3.96×10-4 0.157 0.1180
+Actual Track (Whistler) 1278 149.3 9.76×10-4 0.204 0.1176
+Table 8: Comparison of geometric parameters of the generated and actual Cesana track under different height-difference weighting factors.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+b=0.5 1356.3 122.8 1.60×10-4 0.1237 0.0922
+b=1 1356.7 123.4 1.62×10-4 0.1252 0.0926
+b=1.5 1356.4 123.7 1.63×10-4 0.1265 0.0927
+Actual Track (Cesana) 1359.2 124.5 5.04×10-4 0.183 0.0910
+Table 9: Comparison of geometric parameters of the generated and actual Nagano track under different height-difference weighting factors.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+b=0.5 1702.6 126.4 2.02×10-4 0.0875 0.0744
+b=1 1699.7 126.5 2.04×10-4 0.0878 0.0746
+b=1.5 1695.4 126.7 2.06×10-4 0.0883 0.0749
+Actual Track (Nagano) 1700 118 - 0.150 0.0700
+Table 10: Comparison of geometric parameters of the generated and actual Pyeongchang track under different height-difference weighting factors.
+Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
+b=0.5 1664.5 178.7 1.81×10-4 0.1172 0.1075
+b=1 1664.1 177.8 1.81×10-4 0.1172 0.1074
+b=1.5 1659.7 177.6 1.82×10-4 0.1174 0.1076
+Actual Track (Pyeongchang) 1659 154 - 0.25 0.0948